### v2

# Five-loop Anomalous Dimensions of Cubic Scalar Theory from Operator Product Expansion

Abstract
Recently, an Operator Product Expansion method has been proposed for renormalization in quantum field theories. This method extracts UV divergences from two-point propagator-type integrals, and it is believed to be applicable to generic QFTs. In the present work, we apply the OPE method to the renormalization of scalar theory with cubic interactions in six-dimensions. We have successfully computed the 5-loop corrections to the anomalous dimensions of the $\phi^{Q}$ operator, as well as the large $N$ expansion of scaling dimensions at the Wilson-Fisher fixed point to $1/N^{5}$ order. This computation sets a new loop-order record for the anomalous dimension of the $\phi^{Q}$ operator in cubic scalar theory, and further demonstrates the efficiency and capacity of the OPE method in renormalization.
Keywords:
Anomalous Dimension, OPE, Cubic Scalar Theory, Renormalization
1 Introduction
During the development of the renormalization group method, scalar field theories have served as an ideal playground for testing multi-loop computation techniques and providing insights into more complicated field theories such as non-Abelian gauge theories. Yet the renormalization of scalar field theories is itself of great importance, as explored in Wilson’s theory of critical phenomena Wilson:1971dc , that in certain statistical systems, the critical exponents near second-order phase transitions are related to the anomalous dimensions of the field or composite operators constructed from fields PELISSETTO2002549 ; Moshe:2003xn . Two notable scalar field theories are the six-dimensional cubic scalar field theory and the four-dimensional quartic scalar theory, both of which are renormalizable theories. These two theories belong to the same $O(N)$ universality class, and it has been conjectured that the $O(N)$ quartic scalar theory at its UV fixed point is equivalent to the $O(N)$ cubic scalar theory at its IR fixed point Fei:2014yja ; Fei:2014xta , namely they approach to the same critical theory in the fixed points. Verifying this conjecture with high-loop precision for the scaling dimensions of fields or composite operators represents an intriguing problem, but it demands perturbative computation of renormalization functions to higher-loop orders. Because of the technical advancements, recently there has been a growing interest in pushing renormalization computations to the next loop order.
It has been a long history for renormalization computations of quartic and cubic scalar theories, dating back to the early era studies of critical exponents up to $\epsilon^{2}$ orders for generalized Ising and Baxter models Wilson:1971dc , as well as the generalized Heisenberg model Wilson:1971vs . The 3-loop computations of the beta functions, field renormalization and mass renormalization of the quartic scalar theory were completed just one year after the 1-loop and 2-loop results BREZIN1973227 ; Brezin:1973igb , with 4-loop results following five years later Kazakov:1979ik . The full 5-loop renormalization of the quartic scalar theory required over a decade of collaborative efforts from numerous researchers Chetyrkin:1981jq ; Gorishnii:1983gp ; Kazakov:1983dyk ; Gorishnii:1983jr ; Kleinert:1991rg ; Adzhemyan:2013jra . Throughout this development, techniques such as the Gegenbauer polynomial $x$-space, IR-rearrangement Vladimirov:1979zm ; Chetyrkin:1980pr ; Caswell:1981ek , $R$-operation ($R^{\ast}$-operation) Chetyrkin:1982nn ; CHETYRKIN1984419 ; Larin:2002sc , as well as the modern Feynman diagram generating and Feynman integral reduction algorithms were refined and played the major roles. Early attempts were also made to push renormalization computations to higher-loop order, for instance the primitive divergences of the 7-loop beta function Broadhurst:1995km , or extend them to composite operators, for instance the 4-loop anomalous dimension of operators with derivatives Derkachov:1997pf . However, the complete 6-loop computation of the quartic scalar theory became feasible only approximately 15 years after the 5-loop result, firstly with the 6-loop field renormalization result Batkovich:2016jus , followed by the beta function and mass renormalization for the $O(1)$ model Kompaniets:2016hct , and subsequently for the $O(N)$ model Kompaniets:2017yct . A major breakthrough came with the development of the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which enabled the successful computation of the 7-loop beta function, mass renormalization, and field renormalization Schnetz:2016fhy , and even the 8-loop field renormalization function Schnetz:2025opm . The history of renormalization computations for the cubic scalar theory has closely paralleled that of the quartic scalar theory, which is not surprising since both have benefited from the same technical advancements. However, due to greater complexity, progress in the cubic scalar theory has consistently lagged by 1 or 2 loops compared to its quartic counterpart at similar stages. Early 1-loop and 2-loop analysis of the cubic scalar theory revealed its asymptotic freedom Macfarlane:1974vp ; Ma:1975vn . Over the subsequent 40 years, 3-loop deAlcantaraBonfim:1980pe ; deAlcantaraBonfim:1981sy ; Fei:2014xta and 4-loop Gracey:2015tta renormalization computations were gradually completed, with the 5-loop result finished only very recently Kompaniets:2021hwg ; Borinsky:2021jdb . The frontier record once again came from the graphical function method, by which the 6-loop beta function, field renormalization and mass renormalization functions of the cubic scalar theory have been successfully acquired this very year Schnetz:2025wtu .
Besides the computations of the beta functions and anomalous dimensions of fields and mass, the renormalization of composite operators constitutes an equal important yet more challenging problem. Phenomenologically, this is relevant to specific critical exponents in phase transition problems. One class of composite operators of interest is the fixed-charge symmetric traceless operator $\phi^{Q}$. In statistical physics, the anomalous dimensions of $\phi^{Q}$ for certain values of $Q$ are relevant to phase transitions in diverse systems PELISSETTO2002549 , such as liquid crystals, graphite-intercalation compounds, polymers, etc. DePrato:2003yd ; Calabrese:2004ca . Interest in computing the scaling dimensions of the $\phi^{Q}$ operator in cubic and quartic scalar theories grows from 2019, when the semi-classical method is introduced to compute the scaling dimension of $\phi^{Q}$ for generic $Q$ at the large $Q$ limit. This leads to all-loop results at leading and sub-leading orders in the $Q$ expansion for the quartic scalar theory Arias-Tamargo:2019xld ; Badel:2019oxl ; Antipin:2020abu ; Giombi:2020enj and the cubic scalar theory Arias-Tamargo:2020fow ; Antipin:2021jiw . Progress in the semi-classical approach stimulates the growing attention to the perturbative method approach towards renormalization computation of the $\phi^{Q}$ operator, aiming to validate the semi-classical results and also to obtain results for all orders of $Q$. For the quartic scalar theory, the scaling dimension of the $\phi^{Q}$ operator has been computed perturbatively at 4-loop Jack:2021ypd and 5-loop order Jin:2022nqq . While the 6-loop result is also available Bednyakov:2022guj , in which the data of $Q=1,\ldots,5$ was computed via perturbative Feynman diagram method and the data of $Q=6,7$ was imported from the semi-classical method. Concurrently, for the cubic scalar theory, perturbative computations have been carried out at 3-loop Jack:2021ziq and 4-loop order Huang:2024hsn .
High-loop order renormalization computations would never be possible without the continuous evolution of both traditional and modern techniques. One of the important techniques to mention is the graphical function method Schnetz:2013hqa ; Golz:2015rea ; Borinsky:2021gkd ; Schnetz:2024qqt ; HP , which takes the advantage of computing scalar integrals in coordinate space. It pushes renormalization computations to 8-loops for the four-dimensional quartic scalar theory and 6-loops for the six-dimensional cubic scalar theory, and for the first time, with some tricks it also pushes the renormalization computations to 7-loops in the Gross-Neveu model involving spinning particles Huang:2025ree . The graphical function method is designed to evaluate scalar integrals with no more than four external points, which is certainly not the case for integrals of correlation functions involving $\phi^{Q}$ operators. To address this, we have proposed an Operator Product Expansion (OPE) method for renormalization computations by analyzing the UV divergence structures of correlation functions in the large momentum expansion limit Huang:2024hsn . This OPE method has proven highly useful and efficient for determining anomalous dimensions Huang:2024hsn , and this fact has also been shown in the condensate calculations Marino:2024uco ; Liu:2025bqq . In the OPE framework, the momenta of the two operators are taken to the infinite limit. This can be considered pictorially as two hard fields propagating in a soft field background Huang:2016bmv ; Huang:2022qnx , reducing the integral evaluation to a single-scale problem characterized by the large momentum parameter. By identifying the contributing terms in the OPE, the essential UV divergences required for renormalization computations are found to be determined by two-point hard sub-graphs. Concurrently, the Wilson coefficients in the OPE are proportional to a combination of $Z$-factors and the contributions of these two-point hard sub-graphs. By requiring the UV finiteness of the Wilson coefficients, the $Z$-factors can be determined once the UV divergences of the hard graphs are known. This approach thus reduces renormalization computations to the evaluation of two-point propagator-type integrals. Note that the generality of OPE across any quantum field theories enables diverse applications of this OPE method, to theories with scalars or particles with spins, and also to the situations involving operator mixing and operators with derivatives, with Lorentz indices, etc.
In this work, we apply the OPE method developed in Huang:2024hsn to compute the 5-loop scaling dimension of the $\phi^{Q}$ operators in the six-dimensional $O(N)$ cubic scalar theory, thereby extending beyond the 4-loop result reported in that paper. This paper is structured as follows. In §2, we provide a concise review of the cubic scalar theory, its renormalization functions, and the OPE method for renormalization computation. Treatments for organizing and simplifying Feynman graphs of the correlation functions involving the $\phi^{Q}$ operator are also presented. The §3 is devoted to a detailed demonstration of the OPE method using 2-loop examples, showing how the field renormalization function, beta function and the anomalous dimensions of the $\phi^{Q}$ operator can be computed efficiently. In §4, general discussion for the computation of 5-loop anomalous dimension of $\phi^{Q}$ operator has been presented, and various 5-loop results can be found in this section and the Appendix. In §5, conclusion and discussion are presented.
2 Review of the theory and the method
In this section we review the six-dimensional $O(N)$ cubic scalar theory, with emphasis on its renormalization computations and the techniques for organizing and simplifying Feynman diagrams. To make this paper self-contained, we also provide a brief review of the OPE method for renormalization computations Huang:2024hsn , especially for its application in the context of cubic scalar theory.
2.1 The cubic scalar theory and its renormalization
The theory under consideration is the six-dimensional $O(N)$ cubic scalar theory, defined by the Lagrangian,
$\mathcal{L}=\frac{1}{2}(\partial\phi_{i})^{2}+\frac{1}{2}(\partial\sigma)^{2}-\frac{g}{2!}\sigma\phi_{i}^{2}-\frac{h}{3!}\sigma^{3}~,$ (1)
where $\sigma$ and $\phi_{i},i=1,\ldots,N$ denote two types of real scalar fields, with the $\phi_{i}$ fields transforming as an $O(N)$ vector. This theory is perturbatively renormalizable in six-dimensional space-time. The primary focus of this paper is the renormalization analysis of the symmetric traceless $\phi^{Q}$ operator, defined as,
$\phi^{Q}:=T_{i_{1}\cdots i_{Q}}\phi^{i_{1}}\cdots\phi^{i_{Q}}~,$ (2)
where $T_{i_{1}\cdots i_{Q}}$ is a symmetric traceless tensor and $Q$ is a fixed charge number. We will work in the $\overline{\mbox{MS}}$ regularization scheme tHooft:1973mfk in $d=(6-\epsilon)$-dimension. The classical scaling dimension of $\phi^{Q}$ is Q(d/21)Q(d/2-1), while the total scaling dimension is
$\Delta_{Q}=Q\left(\frac{d}{2}-1\right)+\gamma_{Q}+Q\gamma_{\phi}~.$ (3)
The difference between the total scaling dimension and the classical scaling dimension is defined as the anomalous dimension of the $\phi^{Q}$ operator. The gamma function $\gamma_{Q}$ of $\phi^{Q}$ operator can be computed perturbatively by Feynman diagrams of inserting the $\phi^{Q}$ operator in the 1PI Green function with $Q$ external legs. However, explicit evaluation of these correlation functions is complicated by the cumbersome $O(N)$ tensor structures inherent to the operator, which cause the number of terms in Feynman rules to grow exponentially with increasing $Q$. To address this problem, it is convenient to introduce an auxiliary $O(N)$ zero-norm complex vector $q$ satisfying $^{2}=0, qq$$\cdot\bar{q}=1$, and define the tensor $T$ as Antipin:2020abu ; Jin:2022nqq
$T_{i_{1}\ldots i_{Q}}:=\frac{1}{Q!}q_{i_{1}}q_{i_{2}}\cdots q_{i_{Q}}~.$ (4)
This allows us to redefine the $\phi^{Q}$ operator in terms of a $\varphi^{Q}$ operator,
$\phi^{Q}:=\frac{1}{Q!}\varphi^{Q}~,$ (5)
where the complex field $\varphi$ is a redefinition of the scalar field $\varphi:=q\cdot\phi$, with its complex conjugate field $\bar{\varphi}:=\bar{q}\cdot\phi$. This definition effectively strips $O(N)$ indices from the correlation functions, while the anomalous dimension of $\phi^{Q}$ operator is equivalent to that of $\varphi^{Q}$ operator. The latter can now be computed by $\varphi^{Q}$-$Q\varphi$ correlation functions, which are free of $O(N)$ index structures. In the $\varphi$-field representation, the $\sigma\phi^{2}$ vertices along the line of the operator and the external legs are reformulated as $\sigma\varphi\bar{\varphi}$ vertices, while other $\sigma\phi^{2}$ vertices appearing in loop structures of $\varphi$ fields contribute factors of $N$ after summing over color indices. Due to charge conservation, scalar charge flows along $\varphi$ fields in the corresponding Feynman diagrams.
For purpose of renormalization computation, following standard procedures, we introduce renormalized fields as,
$\varphi_{0}=Z^{\frac{1}{2}}_{\varphi}\varphi~~~,~~~\sigma_{0}=Z^{\frac{1}{2}}_{\sigma}\sigma~,$ (6)
and renormalized coupling constants as,
$g_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{g}g~~~,~~~h_{0}=\mu^{\frac{\epsilon}{2}}\widetilde{Z}_{h}h~~~\mbox{with}~~~\widetilde{Z}_{g}=Z_{g}Z^{-\frac{1}{2}}_{\sigma}Z^{-1}_{\varphi}~~~,~~~\widetilde{Z}_{h}=Z_{h}Z^{-\frac{3}{2}}_{\sigma}~,$ (7)
where $g_{0},h_{0}$ denote bare coupling constants and $\varphi_{0},\sigma_{0}$ denote bare fields. We also define the renormalized operator $\varphi^{Q}_{R}=Z_{\varphi^{Q}}\varphi_{0}^{Q}$. Since $\varphi_{0}^{Q}=Z_{\varphi}^{\frac{Q}{2}}\varphi^{Q}$, we get
$\varphi_{R}^{Q}=\widetilde{Z}_{\varphi^{Q}}\varphi^{Q}~~~\mbox{with}~~~\widetilde{Z}_{\varphi^{Q}}=Z_{\varphi^{Q}}Z_{\varphi}^{\frac{Q}{2}}~.$ (8)
Then $\widetilde{Z}_{\varphi^{Q}}$ can be assigned as the Feynman rule of the $\varphi_{R}^{Q}$-$Q\varphi$ vertex, and the anomalous dimension of the $\varphi^{Q}$ operator can be determined by the $Z_{\varphi^{Q}}$-factor.
Provided that the $Z_{\varphi}$, $Z_{\sigma}$, $Z_{g}$ and $Z_{h}$ have been determined, we can compute beta functions via the standard field theory method. The beta functions $\beta(g)=\frac{\partial g}{\partial\ln\mu}$ and $\beta(h)=\frac{\partial h}{\partial\ln\mu}$ for the renormalized couplings can be derived from $\widetilde{Z}_{h},\widetilde{Z}_{g}$ by solving the following matrix,
$\begin{pmatrix}\frac{1}{g}+\frac{\partial\ln\widetilde{Z}_{g}}{\partial g}&\frac{\partial\ln\widetilde{Z}_{g}}{\partial h}\\ \frac{\partial\ln\widetilde{Z}_{h}}{\partial g}&\frac{1}{h}+\frac{\partial\ln\widetilde{Z}_{h}}{\partial h}\\ \end{pmatrix}\begin{pmatrix}\frac{\partial g}{\partial\ln\mu}\\ \frac{\partial h}{\partial\ln\mu}\\ \end{pmatrix}=\begin{pmatrix}-\frac{\epsilon}{2}\\ -\frac{\epsilon}{2}\\ \end{pmatrix}~$. (9)
Then the field anomalous dimensions of $\varphi$ and $\sigma$ can be computed by definition as,
$\gamma_{\varphi}=\frac{1}{2}\left(\frac{\partial\ln Z_{\varphi}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\varphi}}{\partial h}\beta(h)\right)~~~,~~~\gamma_{\sigma}=\frac{1}{2}\left(\frac{\partial\ln Z_{\sigma}}{\partial g}\beta(g)+\frac{\partial\ln Z_{\sigma}}{\partial h}\beta(h)\right)~.$ (10)
Furthermore, once $Z_{\varphi^{Q}}$ has been determined, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by
$\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}=-\frac{\partial\ln Z_{\varphi^{Q}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}Z_{\varphi}^{-\frac{Q}{2}}}{\partial\ln\mu}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}+Q\left(\frac{1}{2}\frac{\partial\ln Z_{\varphi}}{\partial\ln\mu}\right)~,$ (11)
where the first term computes $\gamma_{Q}:=\gamma_{\varphi^{Q}}$ and the second term computes $Q\gamma_{\varphi}$.
With the knowledge of beta functions and the anomalous dimensions, in order to study the large $N$ expansion of the anomalous dimensions around the Wilson-Fisher fixed point, one need to solve the fixed points from beta functions,
$\beta(g^{\ast})=0~~~,~~~\beta(h^{\ast})=0~.$ (12)
After substituting the solutions $g^{\ast},h^{\ast}$ into the anomalous dimensions and expanding the results in the series of $1/N$, one produces the large $N$ expansion of the cubic scalar theory at the IR fixed point.
The renormalization computation problems require the determination of the $Z$-factors. In general, a $Z$-factor of up to $L$-loop contributions takes the schematic form as
$Z_{{\tiny\texttt{X}}}=1+\sum_{n_{\ell}=1}^{L}\sum_{n_{\epsilon}=1}^{n_{\ell}}\frac{z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}}{\epsilon^{n_{\epsilon}}}~,$ (13)
in which the symbol X can represent the fields $\varphi,\sigma$, the couplings $h,g$ or the operator $\varphi^{Q}$ of arbitrary value $Q$. The $Z_{{\tiny\texttt{X}}}$-factor can be obtained by determining the coefficients $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ up to the appropriate loop orders. All such $Z$-factors can be conveniently computed using two-point propagator-type integrals, following the spirit of the OPE method to renormalization computations. We will review the OPE method in the next subsection.
2.2 The OPE method for the renormalization
In general quantum field theory, the product of two time-ordered operators admits an asymptotic expansion in the limit where their coordinates coincide, known as the Operator Product Expansion. The expansion basis is a set of composite operators of the theory, ordered according to their dimensions, while the expansion coefficients are functions of the coordinate difference, which are in general singular. When considering the OPE in momentum space by Fourier transformation with respect to the coordinate difference, the expansion is performed in the limit of large momentum $p\to\infty$, schematically expressed as,
$\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{ren}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{ren}}(p)~\sim~C_{i}^{\tiny\mbox{ren}}(g_{i};p)\widehat{O}_{i}^{\tiny\mbox{ren}}(k)+\cdots$ (14)
where $k$ is a finite momentum, and the momenta of the two operators tend to the opposite directions of infinite limits [^1] Pictorially, we can consider the two operators with large momentum as two hard fields propagating in the background of soft fields. This idea has also been realized in the study of the boundary operators of BCFW recursion relations Huang:2022qnx . By shifting two external momenta $p_{1},p_{2}$ with a complex auxiliary momentum $zq$ as $p_{1}\to p_{1}+zq$, $p_{2}\to p_{2}-zq$, at the large $z$ limit we get two hard fields propagating in the background of soft fields.. The superscript $^{\tiny\mbox{ren}}$ in the operator $\widehat{O}$ represents a renormalized operator in dimensional regularization scheme.
Alternatively, we may consider the OPE of bare operators, schematically written as,
$\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{bare}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{bare}}(p)~\sim~C_{i}^{\tiny\mbox{bare}}(g_{i0};p)\widehat{O}_{i}^{\tiny\mbox{bare}}(k)+\cdots$ (15)
where the Wilson coefficients are now functions of $p^{2}$ and the bare coupling constants. Since renormalized and bare operators are related by their corresponding $Z$-factor,
$\widehat{O}^{{\tiny\mbox{ren}}}=Z_{\widehat{O}}\widehat{O}^{{\tiny\mbox{bare}}}~,$ (16)
from eqn.(14) and eqn.(15) we get the relation,
$C_{i}^{\tiny\mbox{ren}}(g_{i};p)=\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)~.$ (17)
Special attention should be paid to the OPE of fundamental fields, since the $Z$-factor of a fundamental field $\varphi$ or $\sigma$ does not correspond directly to $Z_{\varphi}$ or $Z_{\sigma}$. Instead, it is defined through $\varphi=Z_{\varphi}^{-\frac{1}{2}}\varphi_{0}$ or $\sigma=Z_{\sigma}^{-\frac{1}{2}}\sigma_{0}$. Furthermore, one usually computes the amputated correlation functions, and an extra factor $Z_{\varphi}$ or $Z_{\sigma}$ should be included for each fundamental field in the OPE of amputated correlation functions. With these two considerations, a fundamental field $\varphi$ or $\sigma$ contributes a $Z^{\frac{1}{2}}_{\varphi}$ or $Z^{\frac{1}{2}}_{\sigma}$ factor to the numerator of the Wilson coefficient relation (17) in the context of amputated OPE. However, the $Z$-factor in the denominator corresponds to an operator but not an external leg, so it still contributes to the $Z^{-\frac{1}{2}}_{\varphi}$ or $Z^{-\frac{1}{2}}_{\sigma}$ factor.
Due to the fact that the Wilson coefficients of renormalized operators $C_{i}^{\tiny\mbox{ren}}(g_{i};p)$ are UV finite, once the Wilson coefficients of bare operators $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed, the ratios of $Z$-factors can be determined using UV finiteness conditions generated from the cancellation of $\epsilon$ poles. It could be possible that some operators in the basis have the same dimension, leading to the operator mixing situation. In this case, the relation (16) is generalized to the vectors of mixed operators related by matrices of $Z$-factors. The $Z$-matrices can be determined by the UV finiteness conditions from all the OPE coefficients corresponding to the mixed operators. Importantly, the above discussion is based on general operators, which means that it is applicable to either simple operators, or operators with Lorentz indices, with fermion and gauge fields, etc.
The Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ of operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ can be computed by the large momentum expansion of bare correlation functions. For an operator $\widehat{O}_{i}^{\tiny\mbox{bare}}$ composed of $n$ fields, the corresponding Wilson coefficient is given by
$C_{i}^{\tiny\mbox{bare}}(g_{i0};p)=\Big{\langle}\widehat{O}_{\tiny\mbox{II}}^{\tiny\mbox{bare}}(k-p)\widehat{O}_{\tiny\mbox{I}}^{\tiny\mbox{bare}}(p)\varphi_{0}(k_{1})\varphi_{0}(k_{2})\cdots\varphi_{0}(k_{n})\Big{\rangle}\Big{|}_{k_{i}\to 0~,~i=1,\ldots,n}~,$ (18)
where $k=-\sum_{i=1}^{n}k_{i}$. Pictorially, this corresponding to splitting the Feynman graph of the original correlation function into hard and soft sub-graphs by cutting $n$ internal propagators and/or external legs, with the Wilson coefficient receiving contributions from all possible $n$-cut hard sub-graphs. If the operator contains fields with derivatives, for instance $\partial_{\mu}\varphi_{0}(k)$, we can compute the coefficient by acting $i\frac{\partial}{\partial k^{\mu}}$ to the bare correlation function. For hard sub-graphs, since all the $k_{i},i=1,\ldots,n$ have been set to zero and there remains only two operators with hard momenta, we effectively get two-point graphs. Hence the Wilson coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ are computed by the two-point propagator-type integrals. This greatly simplifies the computation, especially when dealing with correlation functions with large number of external legs.
The above discussion outlines the core idea of computing renormalization functions via OPE method. To determine beta functions and anomalous dimensions, we first need the $Z$-factors, which are extracted from UV finiteness conditions generated from (17), i.e.,
$\frac{Z_{\widehat{O}_{\tiny\mbox{I}}}Z_{\widehat{O}_{\tiny\mbox{II}}}}{Z_{\widehat{O}_{i}}}C_{i}^{\tiny\mbox{bare}}(g_{i0};p)=\mbox{UV~finite}~.$ (19)
To establish these UV finiteness conditions, we first need to compute the coefficients $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$ from the corresponding two-point integrals. Then using the general ansatz (13) for $Z$-factors, we can expand the complete expression as series expansion of $\epsilon$, and generate equations by setting coefficients of all $\epsilon$ poles to zero. The $z_{\tiny\texttt{X}}^{n_{\ell}n_{\epsilon}}$ coefficients of the $Z_{\tiny\texttt{X}}$-factor can be obtained by solving these equations, and consequently the $Z_{\tiny\texttt{X}}$-factor can be successfully determined to the desired loop order, provided the two-point integrals can be computed to that order.
Now let us apply the OPE method to the renormalization computation of the cubic scalar theory, working in coordinate space. As in the standard field theory computation, the $Z_{\varphi}$ and $Z_{\sigma}$-factors are computed from two-point correlation functions. This computation can be naturally framed within the OPE method. Consider the OPE,
$\varphi\bar{\varphi}~\sim~c_{\mathbf{1}}~\mathbf{1}+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\mathbf{1}}~\mathbf{1}+\cdots~,$ (20)
where $\mathbf{1}$ is the unit operator of zero dimension, and $c$ generically represents Wilson coefficient, which is in general different between the two relations. Since they are the OPE of two fundamental fields, the large momentum expansion of them produces,
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\rangle~\cong~Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))~~~,~~~\langle\sigma(x_{1})\sigma(x_{2})\rangle~\cong~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))~,$ (21)
where we use notation $G(g_{i0};\varphi(x_{1}),\varphi(x_{2}),\varphi(\mbox{\textst{$x$}}_{3}),\ldots)$ to represent the bare Wilson coefficient $C_{i}^{\tiny\mbox{bare}}(g_{i0};p)$. It is computed by the two-point integrals, derived from the bare correlation function with $x$ denoting the corresponding external leg removed, as in eqn.(18). Since they are already two-point functions, no additional legs need removal, so $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}))$ are computed directly from Feynman graphs of two-point correlation functions. Then from UV finiteness conditions of (21), the $Z$-factors of fields $\varphi$ and $\sigma$ can be determined.
The $Z$-factors of coupling constants $g,h$ can be computed from the OPE,
$\varphi\bar{\varphi}~\sim~c_{\sigma}~\sigma+\cdots~~~,~~~\sigma\sigma~\sim~c^{\prime}_{\sigma}~\sigma+\cdots~.$ (22)
In the large momentum expansion, it produces $Z$-factor ratios $Z^{\frac{1}{2}}_{\varphi}Z^{\frac{1}{2}}_{\varphi}/Z^{-\frac{1}{2}}_{\sigma}$ and $Z^{\frac{1}{2}}_{\sigma}Z^{\frac{1}{2}}_{\sigma}/Z^{-\frac{1}{2}}_{\sigma}$ respectively. So we obtain,
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~Z_{\sigma}^{\frac{1}{2}}Z_{\varphi}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (23)
$\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~Z^{\frac{3}{2}}_{\sigma}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (24)
Further using the relation (7) we get [^2] In the previous paper Huang:2024hsn , we have taken the convention that the $g_{0}$ or $h_{0}$ in the denominator is absorbed into the bare correlation function for three-point functions. This leads to slightly different expressions in that paper, where $G$ represents the bare correlation function divided by $g_{0}$ or $h_{0}$ respectively.
$\langle\varphi(x_{1})\bar{\varphi}(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~,$ (25)
$\langle\sigma(x_{1})\sigma(x_{2})\sigma(x_{3})\rangle~\cong~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))~.$ (26)
The $G(g_{0},h_{0};\varphi(x_{1}),\bar{\varphi}(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ and $G(g_{0},h_{0};\sigma(x_{1}),\sigma(x_{2}),\sigma(\mbox{\textst{$x$}}_{3}))$ are directly computed from the two-point integrals of corresponding three-point functions by removing one leg of $\sigma$ field, and the $Z_{g},Z_{h}$ can be determined from UV finiteness conditions of (25), (26). With results of $Z_{\varphi},Z_{\sigma}$ and $Z_{g},Z_{h}$, the beta functions and anomalous dimensions of fields can be readily computed by definitions (9), (10).
The $Z$-factor of the $\varphi^{Q}$ operator can be computed from the OPE,
$\varphi^{Q}\bar{\varphi}~\sim~c_{i}~\varphi^{Q-1}~.$ (27)
After large momentum expansion it produces the ratio of $Z$-factors as,
$\frac{Z_{\varphi^{Q}}Z^{\frac{1}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}}=\frac{Z_{\varphi^{Q}}Z^{\frac{Q}{2}}_{\varphi}}{Z_{\varphi^{Q-1}}Z^{\frac{Q-1}{2}}_{\varphi}}=\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}~.$ (28)
Thus, the $Z_{\varphi^{Q}}$-factor can be computed recursively using the UV finiteness conditions,
$\frac{\widetilde{Z}_{\varphi^{Q}}}{\widetilde{Z}_{\varphi^{Q-1}}}G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mbox{UV~finite}~,$ (29)
where $G_{Q}$ is computed from Feynman graphs of the $\varphi^{Q}$-$Q\varphi$ correlation function by removing (Q1)(Q-1) external legs. They are two-point integrals with large momentum localized at the $\varphi^{Q}$ operator and one remaining external leg. Using the results of $Z_{\varphi^{Q}}$ and the previously computed beta functions, the anomalous dimension of the $\varphi^{Q}$ operator can be computed by definition (11).
2.3 From Feynman diagrams to two-point integrals
The UV finiteness conditions discussed in the previous subsection require computing two-point integrals derived from two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions. In this subsection, we review some techniques which will help to simplify the problem of generating Feynman diagrams and transfer them to two-point graphs for subsequent evaluation of two-point integrals.
The $\Phi$-field representation
The cubic scalar theory we consider involves two types of fields $\varphi,\sigma$, and two types of cubic interactions. The complex field $\varphi$ carries a scalar charge that flows along the line of $\varphi,\bar{\varphi}$ fields, which can be represented by an arrow indicating the direction of charge flow. When generating Feynman diagrams, we need to account for all possible field configurations in the internal lines, which becomes a rather difficult task when the number of internal lines is large. It has also been observed that for a given external field configuration, different internal field configurations only contribute to varying combinations of couplings, while the underlying integrals remain identical. Thus there is no need to generate all Feynman diagrams and redundantly compute the same integrals. To clarify the separation between the structure of couplings and the integrals, we will adopt the $\Phi$-field representation. In this framework, we combine the $\varphi_{i},i=1,\ldots,N$ fields and the $\sigma$ field into a single $(N+1)$ component scalar field as,
$\Phi_{I}=(\Phi_{0},\Phi_{1},\ldots,\Phi_{N}):=(\sigma,\varphi_{1},\ldots,\varphi_{N})~,$ (30)
and the cubic interactions can be universally encoded as
$\frac{1}{6}\lambda_{IJK}\Phi_{I}\Phi_{J}\Phi_{K}~~~\mbox{where}~~~\lambda_{000}=h~~\mbox{and}~~\lambda_{0ij}=g\delta_{ij}~~\forall i,j\neq 0~,$ (31)
with all other components of the symmetric coupling $\lambda_{IJK}$ vanishing. In the $\Phi$-field representation, we essentially generate the topologies of Feynman diagrams without specifying the explicit configuration of component fields. In this case, we have separated the information of the scalar integral part from the factor of couplings. The Feynman scalar integral can be derived from a given topology by assigning propagators according to Feynman rules, while the factors of couplings of all Feynman diagrams sharing the same topology are encoded in the combination of $\lambda_{IJK}$ couplings, which we define as the coupling factor.
The symmetry factor and the coupling factor
When writing the Feynman integral of correlation functions, we first generate all corresponding topologies with private Mathematica code. For each topology, we get the scalar integral from the combination of propagators, multiplied by the standard symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ from the total symmetry of the graph and external legs. Each topology effectively represents a collection of numerous Feynman diagrams that share the same external state configuration but differ in their internal state configurations. By summing over the indices of the $\lambda_{IJK}$, we get the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of the contributing Feynman graphs.
To clarify the concept of the coupling factor, let us take the two-point two-loop topologies $\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ depicted in Fig.(2) as an example.
$Ng^{4}$$Ng^{3}h$$Ng^{3}h$$h^{4}$$g^{3}h$$g^{3}h$$g^{4}$$g^{4}$
Figure 1: Examples of coupling factors in the $\Phi$-field representations and their corresponding component field representations. The internal $\Phi$ fields represent all possible internal state configurations in component fields.
For this topology, we can consider the external field configuration of two external $\sigma$ field represented by the dashed lines, or the field configuration of two external $\varphi,\bar{\varphi}$ fields represented by the arrowed lines, as depicted in the LHS of Fig.(1). Each vertex is assigned a $\lambda_{IJK}$, and after summing over the internal indices, we get the respective coupling factors as,
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\sigma\sigma})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{0I_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}0}=Ng^{4}+2Ng^{3}h+h^{4}~,$ (32)
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\varphi\varphi})=\sum_{I_{1},I_{2},I_{3},I_{4},I_{5}}\lambda_{iI_{1}I_{2}}\lambda_{I_{1}I_{4}I_{3}}\lambda_{I_{2}I_{3}I_{5}}\lambda_{I_{5}I_{4}i}=2g^{3}h+2g^{4}~.$ (33)
Alternatively, we can draw all contributing Feynman diagrams with the explicit internal field configurations, as shown in the RHS of Fig.(1). It can be verified that summing over all four Feynman diagrams of each external field configuration indeed produces the coupling factors evaluated above. Since generating topologies is significantly more efficient than generating complete Feynman diagrams, and the evaluation of coupling factors involves only algebraic operations, the $\Phi$-field representation substantially enhances the computational capability.
The Feynman graphs of $\varphi^{Q}$-$Q\varphi$ correlation functions
In the UV finiteness conditions, the function $G$ is computed schematically as
$G=\sum\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{{\tiny\mbox{topo}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})~,$ (34)
where the summation runs over all possible topologies, and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{{\tiny\mbox{topo}}})$ is computed from the two-point integrals generated by these topologies. Remark that the coupling factor $\mathbb{C}_{\footnotesize\mbox{f}}$ of a topology depends on the explicit external field configuration, but the symmetry factor $\mathbb{S}_{\footnotesize\mbox{ym}}$ and the UV divergence $\mathbb{U}_{\footnotesize\mbox{v}}$ are irrelevant to the external fields, they are computed from the topologies. For two-point correlation functions, $G$ is directly evaluated using all 1PI two-point topologies. For three-point correlation functions, we start from all 1PI three-point topologies and compute the independent two-point graphs by removing one external leg from these topologies in all possible ways. For the $\varphi^{Q}$-$Q\varphi$ correlation function, we start from all 1PI $\varphi^{Q}\to Q\varphi$ topologies, and $G_{Q}$ can be computed from all independent two-point graphs generated by removing (Q1)(Q-1) external legs from these topologies in all possible ways.
By investigating the structure of 1PI $\varphi^{Q}\to Q\varphi$ topologies, it has been proven that only a part of these topologies contributes to the two-point integrals Huang:2024hsn . These contributing topologies are called irreducible graphs, characterized by the property that when the operator vertex is deleted, all remaining external legs stay connected in a single graph (except those that reduce to a single line). The other topologies are called reducible graphs, with the property that when deleting the operator vertex causes the external legs to distribute into two or more separate connected components. Hence we only need to generate the irreducible graphs of the 1PI $\varphi^{Q}\to Q\varphi$ topologies, and compute $G_{Q}$ from their corresponding two-point graphs after all possible removals of (Q1)(Q-1) legs. The summation of these graphs gives,
$G_{Q}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\mathbb{U}_{\footnotesize\mbox{v}}\left(\frac{1}{x_{Q}^{2\lambda}}\right)+\sum_{n=2}^{Q}C_{Q-1}^{n-1}\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})~.$ (35)
The first term comes from the tree-level contribution and evaluates to 1, while
$\mathtt{F}_{n}(\mbox{\textst{$x$}}_{\sigma_{1}},\cdots,\mbox{\textst{$x$}}_{\sigma_{n-1}},x_{Q})=(n-1)!\sum_{n_{\ell}=n-1}^{L}G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})~.$ (36)
The $G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}$ receives contributions from all $n_{\ell}$-loop irreducible topologies. As in the standard procedure, it can be evaluated using the symmetry factor, coupling factor and UV divergence as,
$G_{\scriptsize\mbox{irre}}^{n_{\ell}{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{\sigma_{1}},\ldots,\mbox{\textst{$x$}}_{\sigma_{Q-1}},x_{Q})=\sum_{{\tiny\mbox{irreducible}}\atop{\tiny\mbox{topologies}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{n_{\ell}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\cdots\mbox{\textst{{\scriptsize$\varphi$}}}})~,$ (37)
where the summation is over all non-isomorphic graphs generated by removing (Q1)(Q-1) external legs of the irreducible 1PI topologies of the $\varphi^{Q}$-$Q\varphi$ correlation function.
Let us summarize this section. After generating the independent topologies of $\varphi^{Q}$-$Q\varphi$ correlation functions, one computes the symmetry factor and coupling factor of each topology, and the UV divergences of two-point integrals in the dimensional regularization scheme to an appropriate $\epsilon$ order. Then the $G$ and $G_{Q}$ can be readily computed by summing over all two-point graphs. The $Z$-factors could be determined from UV finiteness conditions, and eventually the anomalous dimensions can be computed from these $Z$-factors.
3 Two-loop demonstration
For illustrative purposes, in this section we will present detailed computations of the two-loop anomalous dimensions of the $\varphi^{Q}$ operator. This requires computing the field renormalization $Z_{\varphi}$ and $Z_{\sigma}$-factors, the coupling renormalization $Z_{h}$ and $Z_{g}$-factors and the $\varphi^{Q}$ operator renormalization $Z_{\varphi^{Q}}$-factor. These factors are determined from the corresponding two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions respectively. All results related to Feynman diagrams are computed by private Mathematica code, while the UV divergences of two-point integrals are computed by HyperlogProcedures Maple package HP . The strategy of computation is formalized as a sequential algorithm, following the computations of the symmetry factors of topologies, the coupling factors of Feynman diagrams, the UV divergence of two-point integrals and finally the determination of $Z$-factors from UV finiteness conditions.
3.1 Two-point correlation functions and the field renormalization
Let us begin with the field renormalization computation. The $Z$-factors $Z_{\varphi}$ and $Z_{\sigma}$ can be computed from the two-point correlation functions $\langle\varphi\bar{\varphi}\rangle$ and $\langle\sigma\sigma\rangle$ respectively as in eqn.(21). The one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(2).
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$
Figure 2: The 1-loop and 2-loop 1PI topologies of two-point functions in $\Phi$-field representation.
Each topology represents a collection of diagrams with distinct allowable internal field configurations of $\sigma,\varphi_{i}$, and these collections are characterized by coupling factors. The symmetry factors and the coupling factors of these topologies are listed in the table below,
topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}})$
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi}$ 1/2 $h_{0}^{2}+Ng_{0}^{2}$ $2g_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi}$ 1/2 $N(2g_{0}^{4}+g_{0}^{2}h_{0}^{2})+h_{0}^{4}$ $Ng_{0}^{4}+2g_{0}^{4}+g_{0}^{2}h_{0}^{2}$
$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi}$ 1/2 $N(g_{0}^{4}+2g_{0}^{3}h_{0})+h_{0}^{4}$ $2g_{0}^{4}+2g_{0}^{3}h_{0}$
There is no need to cut any external legs of these topologies, and their corresponding UV divergences can be directly computed by TwoPoint command in HyperlogProcedures. Feynman integrals in the $\overline{\mbox{MS}}$ dimensional regularization scheme can be evaluated as power series of parameter $\epsilon$ with the results(an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown),
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi})=-\frac{1}{3\epsilon}-\frac{4}{9}+\left(\frac{\pi^{2}}{72}-\frac{13}{27}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (38)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi})=\frac{1}{18\epsilon^{2}}+\frac{43}{216\epsilon}+\left(\frac{1171}{2592}-\frac{\pi^{2}}{216}\right)+\left(-\frac{\zeta_{3}}{108}-\frac{43\pi^{2}}{2592}+\frac{25195}{31104}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (39)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi})=-\frac{1}{3\epsilon^{2}}-\frac{1}{\epsilon}+\left(\frac{\pi^{2}}{36}-\frac{15}{8}\right)+\left(-\frac{7\zeta_{3}}{36}+\frac{\pi^{2}}{12}-\frac{769}{288}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (40)
Then the bare correlation functions are readily obtained as,
$G(g_{0},h_{0};\Phi,\Phi)=1-\sum_{L={\tiny\mbox{L1}},\atop{\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi})~,$ (41)
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factor. Attention should be paid to the $\pm$ signs of individual terms in the summation. The bare correlation function is computed by summing contributions from all tree-level and loop Feynman diagrams. Note that the TwoPoint command only computes the scalar integrals, omitting all $i$ factors from, for instance, the vertices or propagators. When summing these contributions, we must recover the factor $i$ for each loop order. Each vertex carries a factor of i-i. Each propagator contains a factor of $i$ in its numerator, but after Wick rotation to Euclidean space, the denominator acquires an additional factor of 1-1, resulting in an overall factor of i-i for each propagator. Each loop integration gains a factor of $i$ from the Wick rotation $\ell_{0}\to i\ell_{E0}$. Supposing a loop graph has $n_{v}$ vertices, $n_{e}$ propagators, $n_{\ell}$ loops and $n_{p}$ external legs, then the total factor would be $(-i)^{n_{v}}(-i)^{n_{e}}i^{n_{\ell}}$. From graph theory, we have the relations $n_{v}+n_{\ell}-n_{e}=1$ and $3n_{v}=2n_{e}+n_{p}$, and after calculation we get the final factor $(-1)^{n_{p}}i$ for each loop order. For two-point functions, $n_{p}=2$, so each loop order has a factor $i$, whereas the tree-level contrition has a factor i-i. If we factor out an overall i-i, the two-point correlation function can be evaluated as the tree-level contribution minus the sum of loop contributions.
The renormalized two-point correlation functions eqn.(21) should be UV finite, so we get the UV finiteness conditions
$Z_{\varphi}G(g_{0},h_{0};\varphi(x_{0}),\bar{\varphi}(x_{1}))=\mbox{UV~finite}~~~,~~~Z_{\sigma}G(g_{0},h_{0};\sigma(x_{0}),\sigma(x_{1}))=\mbox{UV~finite}~,$ (42)
which means they should be free of $\epsilon$ poles. In order to get the equations of UV finiteness conditions, we rewrite the bare couplings $g_{0},h_{0}$ in the bare correlation functions to $g,h$ by relations eqn.(7), then expand the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$. Note that in the rewriting of couplings with eqn.(7), each $g$ or $h$ is accompanied by a factor $\mu^{\frac{\epsilon}{2}}$, so from the coupling factors we know that at $L$-loop order there will be a $\mu^{L\epsilon}$ factor. Combined with the $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor, at $L$-loop order there will be an overall $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor. Expansion of the factor around $\epsilon\to 0$ produces powers of $\ln\frac{\mu^{2}}{p^{2}}$ terms. Since each $\ln\frac{\mu^{2}}{p^{2}}$ order is independent, the UV finiteness conditions also require the coefficient of each $\ln\frac{\mu^{2}}{p^{2}}$ order is free of $\epsilon$ poles. Hence the higher Log order terms will produce additional equations. Inspecting the UV divergence structures of $L$-loop integrals, the leading divergence is $1/\epsilon^{L}$}, thus at the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ order there will in general be $L$ equations for the cancellation of $L$ poles. At the $\ln^{k}\frac{\mu^{2}}{p^{2}}$ order, the leading divergence is $1/\epsilon^{L-k}$, so there will be (Lk)(L-k) equations. We can either use the equations from the leading Log terms [^3] The leading Log term denotes terms with $\ln^{0}(\mu^{2}/p^{2})$ dependence, and the sub-Log term denotes terms with $\ln(\mu^{2}/p^{2})$ dependence, etc., or include those equations from higher Log terms. If we only use equations of the leading Log terms, then all equations of $Z_{\varphi},Z_{\sigma},Z_{g},Z_{h}$ should be solved simultaneously to completely determine the $Z$-factors. However, if we also use the equations of sub-Log terms, then each $Z$-factor can be determined by solving its own equations of UV finiteness conditions. This will be demonstrated in the following computations.
Now let us take a look at the UV finiteness conditions. For $Z_{\varphi}$-factor, at one-loop, after the above prescribed expansions, we get
$\left(\frac{1}{\epsilon}\left(z_{\phi}^{11}+\frac{g^{2}}{3}\right)+\left(1+\frac{7g^{2}}{9}\right)+\mathcal{O}(\epsilon)\right)+\left(\frac{g^{2}}{3}+\frac{7g^{2}}{9}\epsilon+\mathcal{O}(\epsilon^{2})\right)\ln\frac{\mu^{2}}{p^{2}}+\cdots$ (43)
The coefficients of sub-Log terms are already UV finite, and they will not provide any additional equations. In order for the leading Log term to be finite, the $1/\epsilon$ pole should cancel, and we get
$z_{\varphi}^{11}=-\frac{g^{2}}{3}~.$ (44)
At two-loop order, we need to solve $z^{21}_{\varphi},z^{22}_{\varphi}$. Again after expansions, the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles at leading Log term produces two equations,
$0=z_{\varphi}^{21}+z_{\varphi}^{11}+\frac{7}{3}\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{g^{2}}{3}-\frac{67g^{4}N}{432}+\frac{293g^{4}}{216}+\frac{5g^{3}h}{3}-\frac{67g^{2}h^{2}}{432}\right)~,$
$0=z_{\varphi}^{22}+\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{4}N}{36}-\frac{g^{2}h^{2}}{36}\right)~,$
and the vanishing of $1/\epsilon$ pole at sub-Log term produces one equation,
$0=\left(\frac{2}{3}g^{2}z_{g}^{11}-\frac{1}{3}g^{2}z_{\varphi}^{11}-\frac{1}{3}g^{2}z_{\sigma}^{11}\right)+\left(\frac{5g^{4}}{9}+\frac{2g^{3}h}{3}-\frac{g^{4}N}{18}-\frac{g^{2}h^{2}}{18}\right)~.$
In the traditional approach, the renormalization computation should be performed order by order. This is similar to the situation that solving the $Z$-factors using only equations of leading Log terms. For instance, by using two equations of leading Log term shown above, the determination of $z_{\varphi}^{21},z_{\varphi}^{22}$ will also rely on the results of all one-loop $Z$-factors, including $Z_{g},Z_{h},Z_{\varphi},Z_{\sigma}$, which must be determined a priori. Alternatively, if considering the UV finiteness conditions of all Log terms of a correlation function, the generated equations are sufficient to completely solve the corresponding $Z$-factor. This means that the computation does not rely on other $Z$-factors of lower-loop orders. For instance, the equation from sub-Log term shown above can be solved to obtain a combination of one-loop $Z$-factors. This one-loop $Z$-factor combination is exactly the one appearing in the equations of leading Log term, and after substituting it back to those two equations, we can completely determine the two-loop $Z_{\varphi}$-factor. Furthermore, as a by-product, the equations of sub-Log terms provide constraints to other $Z$-factors at lower-loop orders, which can be served as consistent checks for lower-loop $Z$-factors computed by other correlation functions.
Consequently, we determine the two-loop $Z_{\varphi}$-factors as,
$z_{\varphi}^{21}=\frac{11g^{4}N}{432}-\frac{13g^{4}}{216}-\frac{g^{3}h}{9}+\frac{11g^{2}h^{2}}{432}~~~,~~~z_{\varphi}^{22}=-\frac{g^{4}N}{36}+\frac{5g^{4}}{18}+\frac{g^{3}h}{3}-\frac{g^{2}h^{2}}{36}~.$ (45)
The determination of $Z_{\sigma}$-factor follows exactly the same procedures, and the only difference is the coupling factors. After the same computations, eventually we determine the two-loop $Z_{\sigma}$-factor as
$z_{\sigma}^{11}=-\frac{g^{2}N}{6}-\frac{h^{2}}{6}~,$ (46)
$z_{\sigma}^{21}=\frac{11g^{2}h^{2}N}{432}-\frac{g^{4}N}{216}-\frac{g^{3}hN}{9}-\frac{13h^{4}}{432}~~~,~~~z_{\sigma}^{22}=\frac{g^{4}N}{9}+\frac{g^{3}hN}{3}-\frac{g^{2}h^{2}N}{36}+\frac{5h^{4}}{36}~.$ (47)
The higher-loop $Z_{\varphi},Z_{\sigma}$-factors can be determined following the same procedures. However there would be much more Feynman graphs, and the evaluation of two-point integrals would become considerably more difficult.
3.2 Three-point correlation functions and the beta functions
The determination of $Z_{h},Z_{g}$-factors is quite similar as that of $Z_{\varphi},Z_{\sigma}$-factors. We start from OPE and arrive at the UV finiteness conditions as eqn.(25), (26). Once the large momentum expansion of the bare correlation functions are computed, we can determine $Z_{g}$ and $Z_{h}$ from the UV finiteness conditions. Remind that the UV divergence is computed by all possible three-point topologies with one leg removed. Following the computation strategy, the one-loop and two-loop topologies in $\Phi$-field representation are sketched in Fig.(3), as well as their independent cut graphs.
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-I}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-II}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\Phi}$$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$
Figure 3: The 1-loop and 2-loop 1PI topologies of three-point functions in $\Phi$-field representation. A gray arrow points to the cut graphs of each topology. They are generated by the non-isomorphic computation of all possible removals of one external leg.
The $Z_{g}$-factor is determined by $\langle\varphi\bar{\varphi}\sigma\rangle$ with a $\sigma$ field removed, and the $Z_{h}$-factor is determined by $\langle\sigma\sigma\sigma\rangle$ with a $\sigma$ field removed. The symmetry factors and coupling factors of cut graphs are listed in Table.(1).
topology $\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\sigma\sigma\mbox{\textst{{\scriptsize$\Phi$}}}})$ $\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma_{\varphi\bar{\varphi}\mbox{\textst{{\scriptsize$\Phi$}}}})$
$\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $g^{3}N+h^{3}$ $g^{3}+g^{2}h$
$\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1/2 $2g^{5}N+g^{2}h^{3}N+h^{5}$ $g^{5}N+2g^{4}h+g^{3}h^{2}$
$\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $2g^{5}N+g^{2}h^{3}N+h^{5}$ $2g^{5}+g^{4}hN+g^{2}h^{3}$
$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1 $g^{5}N+g^{4}hN+g^{3}h^{2}N+h^{5}$ $g^{5}N+g^{5}+g^{4}h+g^{2}h^{3}$
$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 2 $g^{5}N+g^{4}hN+g^{3}h^{2}N+h^{5}$ $g^{5}+2g^{4}h+g^{3}h^{2}$
$\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}}$ 1/2 $3g^{4}hN+h^{5}$ $2g^{5}+2g^{3}h^{2}$
Table 1: The symmetry factors and coupling factors of cut graphs for 2-loop three-point topologies.
The UV divergence is computed by TwoPoint command for two-point integral of cut graphs, with the results [^4] Again an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ for $L$-loop integral has not been explicitly shown,
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{\epsilon}+3+\left(\frac{23}{4}-\frac{\pi^{2}}{16}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (48)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2671}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{80911}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (49)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{67}{72\epsilon}-\left(\frac{2743}{864}-\frac{5\pi^{2}}{288}\right)-\left(\frac{86167}{10368}-\frac{335\pi^{2}}{3456}-\frac{5\zeta_{3}}{144}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (50)
and
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{21}{8\epsilon}+\left(\frac{275}{32}-\frac{5\pi^{2}}{96}\right)+\left(\frac{2785}{128}-\frac{35\pi^{2}}{128}-\frac{5\zeta_{3}}{48}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (51)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{21}{8\epsilon}+\left(\frac{259}{32}-\frac{5\pi^{2}}{96}\right)+\left(\frac{7139}{384}-\frac{35\pi^{2}}{128}+\frac{19\zeta_{3}}{48}\right)\epsilon+\mathcal{O}(\epsilon^{2})~,$ (52)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=\frac{1}{2\epsilon}+\left(\frac{83}{24}-\zeta_{3}\right)+\left(\frac{3983}{288}-\frac{5\pi^{2}}{96}-\frac{16\zeta_{3}}{3}-\frac{\pi^{4}}{120}\right)\epsilon+\mathcal{O}(\epsilon^{2})~.$ (53)
Special attention should be paid to the tree-level contribution, which is evaluated by the two-point integral of a three-point vertex with one leg removed. Evaluated in the coordinate space, the result is not 1, but
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L0}}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})=1+\epsilon+\left(\frac{3}{4}-\frac{\pi^{2}}{48}\right)\epsilon^{2}+\left(\frac{1}{2}-\frac{\pi^{2}}{48}-\frac{\zeta_{3}}{24}\right)\epsilon^{3}+\left(\frac{5}{16}-\frac{\pi^{2}}{64}-\frac{\zeta_{3}}{24}+\frac{\pi^{4}}{23040}\right)\epsilon^{4}+\mathcal{O}(\epsilon^{5})~.$ (54)
Appropriate $\epsilon$ order terms should be included for the $L$-loop renormalization computation. Then the bare correlation functions for two-loop computation are readily obtained as [^5] For three-point function, $n_{p}=3$. As discussed in the two-point case, each loop order has a factor i-i, while at the tree-level the factor is also i-i. If factoring out an overall factor i-i, the three-point correlation function can be evaluated by tree contribution plus all loop contributions.,
$G(g_{0},h_{0};\Phi,\Phi,\mbox{\textst{$\Phi$}})=g_{\Phi}(1+\epsilon)+\sum_{L={\tiny\mbox{L1}},{\tiny\mbox{L2-I}}_{1},{\tiny\mbox{L2-I}}_{2},\atop{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},{\tiny\mbox{L2-III}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{L}_{\Phi\Phi\mbox{\textst{{\scriptsize$\Phi$}}}})~,$ (55)
where the field $\Phi$ takes the corresponding component $\sigma$ or $\varphi$ in $\mathbb{C}_{\footnotesize\mbox{f}}$ factors, and $g_{\Phi}=g$ for $\langle\varphi\bar{\varphi}\sigma\rangle$ while $g_{\Phi}=h$ for $\langle\sigma\sigma\sigma\rangle$.
The UV finiteness conditions of eqn.(25) and eqn.(26) can be realized by,
$\frac{Z_{g}g}{g_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\varphi,\bar{\varphi},\mbox{\textst{$\sigma$}})=\mbox{UV finite}~~,~~\frac{Z_{h}h}{h_{0}}\mu^{\frac{\epsilon}{2}}G(g_{0},h_{0};\sigma,\sigma,\mbox{\textst{$\sigma$}})=\mbox{UV~finite}~.$ (56)
After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limit $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get equations for $Z$-factors by the vanishing of $\epsilon$ poles in each $\ln\frac{\mu^{2}}{p^{2}}$ term [^6] The bare coupling in the denominator will introduce a $\mu^{\frac{\epsilon}{2}}$, which cancels the $\mu^{\frac{\epsilon}{2}}$ factor in numerator. In the bare correlation function, at $L$-loop order there will be a $(1/p^{2})^{\frac{\epsilon}{2}L}$ factor from integration, and a $\mu^{\frac{\epsilon}{2}(2L+1)}$ factor from the couplings. If we take out an overall factor $\mu^{\frac{\epsilon}{2}}$ which will not affect the $\epsilon$ expansion, then each $L$-loop is assigned with a $(\mu^{2}/p^{2})^{\frac{\epsilon}{2}L}$ factor contributing to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion at $\epsilon\to 0$ limit..
For $Z_{g}$-factor, at one-loop there is only one equation from the vanishing of $1/\epsilon$ pole of the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term, which yields,
$g^{3}+g^{2}h+gz_{g}^{11}=0~~~\to~~~z_{g}^{11}=-g^{2}-gh~.$ (57)
At two-loop order, there will be two equations from the $\ln^{0}\frac{\mu^{2}}{p^{2}}$ term,
$0=z_{g}^{22}+z_{g}^{21}+(9g^{2}+6gh+1)z_{g}^{11}+3g\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~+\left(\frac{311g^{4}N}{144}+\frac{469g^{4}}{72}-\frac{67g^{3}hN}{72}+\frac{439g^{3}h}{36}+\frac{761g^{2}h^{2}}{144}+g^{2}+\frac{61gh^{3}}{36}+gh\right)~,$ (58)
$0=z_{g}^{22}+(2gh+3g^{2})z_{g}^{11}+g\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}\right)~,$ (59)
and one equation from the sub-Log term,
$0=(3g+2h)z_{g}^{11}+\Big{(}hz_{h}^{11}-(2g+h)z_{\phi}^{11}-(g+2h)z_{\sigma}^{11}\Big{)}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{5g^{3}N}{6}+\frac{7g^{3}}{3}-\frac{g^{2}hN}{3}+\frac{14g^{2}h}{3}+\frac{11gh^{2}}{6}+\frac{2h^{3}}{3}\right)~.$ (60)
Again we confirm that the one-loop $Z$-factors can be solved as a whole by the equation of sub-Log term. Consequently, from the equations of leading Log terms we get
$z_{g}^{21}=\frac{7g^{3}hN}{72}-\frac{11g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{19g^{3}h}{36}-\frac{101g^{2}h^{2}}{144}-\frac{gh^{3}}{36}~~,~~z_{g}^{22}=\frac{5g^{4}N}{12}+\frac{7g^{4}}{6}-\frac{g^{3}hN}{6}+\frac{7g^{3}h}{3}+\frac{11g^{2}h^{2}}{12}+\frac{gh^{3}}{3}~.$
Similarly, solving the equations of UV finiteness conditions for $Z_{h}$-factors, we determine the two-loop $Z_{h}$-factor. We get $z_{h}^{11}=-(g^{3}N+h^{3})/h$, and
$z_{h}^{21}=\frac{7g^{2}h^{2}N}{48}-\frac{g^{5}N}{12h}-\frac{9g^{4}N}{8}-\frac{3g^{3}hN}{8}-\frac{23h^{4}}{48}~~,~~z_{h}^{22}=\frac{g^{5}N}{h}+\frac{3g^{4}N}{2}+\frac{3g^{3}hN}{2}-\frac{g^{2}h^{2}N}{4}+\frac{5h^{4}}{4}~.$
The higher-loop $Z_{h},Z_{g}$-factors can be similarly determined by including more graphs and more orders of $\epsilon$ series of two-point integrals.
With knowledge of $Z_{h},Z_{g},Z_{\phi}$ and $Z_{\sigma}$, we can compute beta functions by solving equations (9). After expanding the solutions at the limits $g,h\to 0$, we get two-loop results as,
$\beta(g)$ $=$ $-\frac{\epsilon}{2}g+\left(\frac{g^{3}N}{12}-\frac{2g^{3}}{3}-g^{2}h+\frac{gh^{2}}{12}\right)$ (61)
$+\left(\frac{11g^{4}hN}{36}+\frac{13gh^{4}}{432}-\frac{43g^{5}N}{216}-\frac{67g^{5}}{54}-\frac{5g^{4}h}{6}-\frac{11g^{3}h^{2}N}{432}-\frac{157g^{3}h^{2}}{108}-\frac{g^{2}h^{3}}{18}\right)~,$~~
$\beta(h)$ $=$ $-\frac{\epsilon}{2}h+\left(\frac{g^{2}hN}{4}-g^{3}N-\frac{3h^{3}}{4}\right)+\left(\frac{31g^{2}h^{3}N}{144}-\frac{g^{5}N}{6}-\frac{161g^{4}hN}{72}-\frac{5g^{3}h^{2}N}{12}-\frac{125h^{5}}{144}\right)~.$~~ (62)
The 2-loop anomalous dimension of field $\varphi$ or $\sigma$ is then given by definition (10) as,
$\gamma_{\varphi}=\frac{g^{2}}{6}+\frac{13g^{4}}{216}+\frac{g^{3}h}{9}-\frac{11g^{4}N}{432}-\frac{11g^{2}h^{2}}{432}~~~,~~~\gamma_{\sigma}=\frac{h^{2}}{12}+\frac{g^{2}N}{12}+\frac{g^{4}N}{216}+\frac{g^{3}hN}{9}-\frac{11g^{2}h^{2}N}{432}+\frac{13h^{4}}{432}~.$ (63)
3.3 Anomalous dimensions of the $\varphi^{2}$ and $\varphi^{3}$ operators
The $\widetilde{Z}_{\varphi^{Q}}$-factors for the $\varphi^{Q}$ operator are determined by the UV finiteness conditions (29), and the corresponding two-point integrals are evaluated by (35). When $Q=2$, all contributing topologies are irreducible topologies. The tree-level, 1-loop and 2-loop topologies of $\varphi^{2}$-$2\varphi$ correlation function have been sketched in Fig.(4), together with their corresponding cut graphs.
$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi}$$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}i}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-II}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-III}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}}$
Figure 4: The tree-level, 1-loop and 2-loop 1PI topologies of $\varphi^{2}$-$2\varphi$ correlation function, where the in subscript denotes a $\phi^{Q}$ operator. A gray arrow points to the cut graphs of each topology. They are generated by the non-isomorphic computation of all possible removals of one external leg. The UV divergences are computed from these cut graphs.
Note that due to the symmetry of graphs, a topology could generate more than one non-isomorphic cut graphs, similar as we have mentioned in the three-point function computation. For the cut graphs related to the same irreducible topology, the symmetry factors as well as the coupling factors are the same as those of the topology. Now Let us compute the $Z$-factor of $\varphi^{2}$, from eqn.(35) we get the explicit expression for $\varphi^{2}$ as
$G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})=1+\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})~.$ (64)
There is only one 1-loop irreducible topology $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi{\varphi}}$, which generates one cut graph $\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}}$. With results,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g^{2}_{0}~~~,~~~\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{\epsilon}+2+\epsilon\left(\frac{5}{2}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (65)
we get the required UV contribution of all 1-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ as
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (66)
Note that in (65), an overall factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ with $L=1$ is not explicitly shown in the result of UV divergence, while there will also be a factor $\mu^{\epsilon L}$ from the couplings $\mathbb{C}_{\footnotesize\mbox{f}}$ at $L$-loop. Hence for (66), if recovering the $(1/p^{2})^{\frac{\epsilon}{2}}$ factor and turning $g_{0},h_{0}\to g,h$, there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{\frac{\epsilon}{2}}$ factor, which contributes to the $\ln\frac{\mu^{2}}{p^{2}}$ expansion. Generally, for $G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}$ there will be an overall $\left(\frac{\mu^{2}}{p^{2}}\right)^{j\frac{\epsilon}{2}}$ factor. For presentation purpose, the factor $(1/p^{2})^{\frac{\epsilon}{2}L}$ will always be treated as implicit in the expressions of UV divergences.
The 1-loop $\widetilde{Z}_{\varphi^{2}}$ can be determined from the UV finiteness of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, where we have used the fact that the 1PI contribution of $\widetilde{Z}_{\varphi^{1}}=1$. After rewriting all bare couplings $g_{0},h_{0}$ to $g,h$ by relations eqn.(7), then expanding the full expression at the limits $g,h\to 0$ followed by the limit $\epsilon\to 0$, we get the condition for the vanishing of $1/\epsilon$ pole in the leading Log term as,
$0=z_{\varphi^{2}}^{11}+g^{2}~~~\to~~~z_{\varphi^{2}}^{11}=-g^{2}~,$ (67)
and the sub-Log term provides no more conditions. Similarly, in order to determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$, the contribution of $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ should be included. There are 5 independent irreducible topologies, which generates 7 cut graphs. The symmetry factors and coupling factors can be simply computed as,
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~,~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~,$ (68)
and
$\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}N+g_{0}^{2}h_{0}^{2}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}+g_{0}^{3}h_{0}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=2g_{0}^{4}~~,~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (69)
The UV divergences are again computed by TwoPoint command in HyperlogProcedures, which are
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1759}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{42487}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (70)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{155}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(\frac{5\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{2915}{384}\right)+\mathcal{O}(\epsilon^{2})~,$ (71)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-II}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1313}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (72)
and
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1687}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{38095}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (73)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-III}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{6\epsilon^{2}}-\frac{55}{72\epsilon}+\left(\frac{\pi^{2}}{72}-\frac{1759}{864}\right)+\epsilon\left(\frac{\zeta_{3}}{36}+\frac{55\pi^{2}}{864}-\frac{42487}{10368}\right)+\mathcal{O}(\epsilon^{2})~,$ (74)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{155}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(\frac{5\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{2915}{384}\right)+\mathcal{O}(\epsilon^{2})~,$ (75)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\left(\frac{71}{24}-\zeta_{3}\right)+\epsilon\left(-\frac{\pi^{4}}{120}-\frac{13\zeta_{3}}{3}-\frac{\pi^{2}}{24}+\frac{2807}{288}\right)+\mathcal{O}(\epsilon^{2})~.$ (76)
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{2}\to 2\varphi$ is
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})=\sum_{L={\tiny\mbox{L2-I}},{\tiny\mbox{L2-II}}_{1},{\tiny\mbox{L2-II}}_{2},\atop{\tiny\mbox{L2-III}}_{1},{\tiny\mbox{L2-III}}_{2},{\tiny\mbox{L2-IX}},{\tiny\mbox{L2-X}}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (77)
After the series expansion of ${\widetilde{Z}_{\varphi^{2}}}G_{Q=2}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{Q})$, we get the UV finiteness conditions for the vanishing of $1/\epsilon$ and $1/\epsilon^{2}$ poles in the leading Log term as,
$0=z_{\varphi^{2}}^{21}+\left(2g^{2}+1\right)z_{\varphi^{2}}^{11}+\left(g^{2}-\frac{7g^{4}N}{144}+\frac{193g^{4}}{72}+\frac{g^{3}h}{4}-\frac{7g^{2}h^{2}}{144}\right)~,$ (78)
$0=z_{\varphi^{2}}^{22}+g^{2}z_{\varphi^{2}}^{11}+\left(\frac{g^{4}N}{12}-\frac{g^{4}}{6}-g^{3}h+\frac{g^{2}h^{2}}{12}\right)~.$ (79)
The $z_{\varphi^{2}}^{11}$ has already been determined in the 1-loop computation, thus by solving above two equations we get,
$z_{\varphi^{2}}^{21}=\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~~~,~~~z_{\varphi^{2}}^{22}=-\frac{g^{4}N}{12}+\frac{7g^{4}}{6}+g^{3}h-\frac{g^{2}h^{2}}{12}~.$ (80)
Note that the vanishing of $1/\epsilon$ pole in the sub-Log term provides an additional equation $0=g^{2}z_{\varphi^{2}}^{11}+g^{4}$. Hence even if we did not compute $z_{\varphi^{2}}^{11}$ at 1-loop, equations of leading Log and sub-Log terms are sufficient to solve $z_{\varphi^{2}}^{11},z_{\varphi^{2}}^{21},z_{\varphi^{2}}^{22}$, and completely determine the 2-loop $\widetilde{Z}_{\varphi^{2}}$-factor.
The determination of $\widetilde{Z}_{\varphi^{3}}$-factor follows exactly the same procedure, where now we should use the UV finiteness conditions of $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$. The topologies of $\varphi^{3}$-$3\varphi$ correlation function are sketched in Fig.(5).
$\Gamma^{\tiny\mbox{L0}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L1}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-I}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-II}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-III}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-IX}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-X}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\varphi\varphi}$$\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{1}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{2}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$$\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\tiny$\varphi$}}}\mbox{\textst{{\tiny$\varphi$}}}}$
Figure 5: The tree-level, 1-loop and 2-loop irreducible 1PI topologies of $\varphi^{3}$-$3\varphi$ correlation function. The topologies in the first row are equivalent to those corresponding to the $\varphi^{2}$-$2\varphi$ correlation function. There are only 2 new irreducible topologies shown in the second row. A gray arrow points to the cut graphs of each new irreducible topology, which are generated by the non-isomorphic computation of all possible removals of two external legs.
Notice that the loop topologies in the first row are equivalent to the $\varphi^{2}\to 2\varphi$ irreducible topologies we have already computed. There are only 2 new irreducible topologies $\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi{\varphi}\varphi}$, $\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi{\varphi}\varphi}$ at 2-loops for $\varphi^{3}\to 3\varphi$. Again from eqn.(35) we get the explicit expression for $\varphi^{3}$ as
$G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})=1+C_{2}^{1}\sum_{j=1}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})+(2!)C_{2}^{2}\sum_{j=2}^{L}G_{\scriptsize\mbox{irre}}^{j{\tiny\mbox{-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})~.$ (81)
Since $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{1-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ and $G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},x_{2})$ have already been computed in $\varphi^{2}$ case as (66) and (77), in order to compute 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor, we only need to compute the additional contributions from these 2 new irreducible topologies of $\varphi^{3}\to 3\varphi$. They generate in total 4 cut graphs as shown in Fig.(5). Their symmetry factors and coupling factors are given by
$\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2}~~~,~~~\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=1~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{3}h_{0}~~~,~~~\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L2-XII}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=g_{0}^{4}~.$ (82)
The UV divergences computed by TwoPoint command are,
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XI}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon}+\frac{17}{8}+\epsilon\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\mathcal{O}(\epsilon^{2})~,$ (83)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{1}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{163}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1161}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (84)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{2}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=-\frac{1}{2\epsilon^{2}}-\frac{15}{8\epsilon}+\left(\frac{\pi^{2}}{24}-\frac{137}{32}\right)+\epsilon\left(\frac{\zeta_{3}}{12}+\frac{5\pi^{2}}{32}-\frac{971}{128}\right)+\mathcal{O}(\epsilon^{2})~,$ (85)
$\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L2-XII}_{3}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})=\frac{1}{2\epsilon^{2}}+\frac{17}{8\epsilon}+\left(\frac{171}{32}-\frac{\pi^{2}}{24}\right)+\epsilon\left(-\frac{\zeta_{3}}{12}-\frac{17\pi^{2}}{96}+\frac{1313}{128}\right)+\mathcal{O}(\epsilon^{2})~.$ (86)
Then the required UV contribution of all 2-loop irreducible topologies for $\varphi^{3}\to 3\varphi$ is
$G_{\scriptsize\mbox{irre}}^{{\tiny\mbox{2-loop}}}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{3})=\sum_{L={\tiny\mbox{L2-XI}},{\tiny\mbox{L2-XII}}_{1}\atop{\tiny\mbox{L2-XII}}_{2},{\tiny\mbox{L2-XII}}_{3}}\mathbb{S}_{\footnotesize\mbox{ym}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{C}_{\footnotesize\mbox{f}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})\mathbb{U}_{\footnotesize\mbox{v}}(\Gamma^{\tiny\mbox{L}}_{\otimes\varphi\mbox{\textst{{\scriptsize$\varphi$}}}\mbox{\textst{{\scriptsize$\varphi$}}}})~.$ (87)
With the previously computed $\widetilde{Z}_{\varphi^{2}}$-factor, we can take the series expansion for $\frac{\widetilde{Z}_{\varphi^{3}}}{\widetilde{Z}_{\varphi^{2}}}G_{Q=3}(g_{0},h_{0};x_{0},\mbox{\textst{$x$}}_{1},\mbox{\textst{$x$}}_{2},x_{Q})$, and get the following equations from the vanishing of $1/\epsilon$, $1/\epsilon^{2}$ poles in the leading Log term and $1/\epsilon$ pole in the sub-Log term,
$0=z_{\varphi^{3}}^{21}+(4g^{2}+1)z_{\varphi^{3}}^{11}+\left(3g^{2}-\frac{7g^{4}N}{48}+\frac{355g^{4}}{24}+\frac{5g^{3}h}{4}-\frac{7g^{2}h^{2}}{48}\right)~,$ (88)
$0=z_{\varphi^{3}}^{22}+3g^{2}z_{\varphi^{3}}^{11}+\left(\frac{g^{4}N}{4}+\frac{5g^{4}}{2}-3g^{3}h+\frac{g^{2}h^{2}}{4}\right)~~~,~~~0=2g^{2}z_{\varphi^{3}}^{11}+6g^{4}~.$ (89)
By solving these 3 equations we completely determine the 2-loop $\widetilde{Z}_{\varphi^{3}}$-factor as
$z_{\varphi^{3}}^{11}=-3g^{2}~~,~~z_{\varphi^{3}}^{21}=\frac{7g^{4}N}{48}-\frac{67g^{4}}{24}-\frac{5g^{3}h}{4}+\frac{7g^{2}h^{2}}{48}~~,~~z_{\varphi^{3}}^{22}=-\frac{g^{4}N}{4}+\frac{13g^{4}}{2}+3g^{3}h-\frac{g^{2}h^{2}}{4}~.$ (90)
With results (67), (80) and (90), the 2-loop $\widetilde{Z}_{\varphi^{2}}$, $\widetilde{Z}_{\varphi^{3}}$-factors are finally determined as,
$\widetilde{Z}_{\varphi^{2}}=1+\frac{z_{\varphi^{2}}^{11}}{\epsilon}+\frac{z_{\varphi^{2}}^{21}}{\epsilon}+\frac{z_{\varphi^{2}}^{22}}{\epsilon^{2}}~~~,~~~\widetilde{Z}_{\varphi^{3}}=1+\frac{z_{\varphi^{3}}^{11}}{\epsilon}+\frac{z_{\varphi^{3}}^{21}}{\epsilon}+\frac{z_{\varphi^{3}}^{22}}{\epsilon^{2}}~.$ (91)
The gamma functions of $\varphi^{Q}$ can be computed by $\gamma_{\varphi^{Q}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{Q}}}{\partial\ln\mu}$, hence we get
$\gamma_{\varphi^{2}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{2}}}{\partial h}\beta(h)=-g^{2}+\left(\frac{7g^{4}N}{72}-\frac{49g^{4}}{36}-\frac{g^{3}h}{2}+\frac{7g^{2}h^{2}}{72}\right)~,$ (92)
$\gamma_{\varphi^{3}}=-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial g}\beta(g)-\frac{\partial\ln\widetilde{Z}_{\varphi^{3}}}{\partial h}\beta(h)=-3g^{2}+\left(\frac{7g^{4}N}{24}-\frac{67g^{4}}{12}-\frac{5g^{3}h}{2}+\frac{7g^{2}h^{2}}{24}\right)~,$ (93)
where $\beta(g),\beta(h)$ have been provided in (61), (62). By the ansatz that the 2-loop $\gamma_{\varphi^{Q}}$ takes the form,
$\gamma_{\varphi^{Q}}=c_{1}Q(Q-1)(Q-2)+c_{2}Q(Q-1)~,$ (94)
and from results $\gamma_{\varphi^{2}},\gamma_{\varphi^{3}}$ we determine
$c_{1}=-\frac{g^{4}}{4}-\frac{g^{3}h}{6}~~~,~~~c_{2}=-\frac{g^{2}}{2}+\frac{7g^{4}N}{144}-\frac{49g^{4}}{72}-\frac{g^{3}h}{4}+\frac{7g^{2}h^{2}}{144}~.$ (95)
Then the scaling dimension of the $\varphi^{Q}$ operator can be calculated by its definition as,
$\Delta_{Q}=\left(\frac{d}{2}-1\right)Q+\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}~,$ (96)
where the anomalous dimension of the field $\varphi$ has been computed in (63).
4 Five-loop anomalous dimensions
The procedures described in previous section for 2-loop computations can be straightforwardly generalized to higher-loop calculations following exactly the same manner, with the primary difference being the need to consider a larger number of graphs. The $Z_{\varphi},Z_{\sigma}$-factors are determined from their corresponding two-point topologies in the manner analogous to (41). Similarly, the $Z_{h},Z_{g}$-factors are determined from the two-point integrals of cut graphs, generated by removing one external leg from the respective three-point topologies, following the approach outlined in (55). The beta functions are then computed from $\widetilde{Z}_{g},\widetilde{Z}_{h}$ using the definition (9). The field anomalous dimensions $\gamma_{\varphi}$ and $\gamma_{\sigma}$ can be computed from the $Z_{\varphi},Z_{\sigma}$-factors combined with the beta functions, following the definition (10). The $\widetilde{Z}_{\varphi^{Q}}$ factor is determined from the two-point integrals of cut graphs generated by removing (Q1)(Q-1) external legs from irreducible 1PI $\varphi^{Q}\to Q\varphi$ topologies, in a manner similar to that described in eqn.(35). The anomalous dimension of the $\varphi^{Q}$ operator is then given by $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$, where $\gamma_{\varphi^{Q}}$ is computed from $\widetilde{Z}_{\varphi^{Q}}$ and the beta functions by its definition. Determining the 5-loop $\gamma_{\varphi^{Q}}$ for general $Q$ requires the results of $L$-loop $\gamma_{\varphi^{Q_{L}}}$ with $Q_{L}=2,\ldots,L+1$ up to 5-loops. Since UV divergences are computed from scalar integrals, we do not need to consider the specific field configurations of Feynman diagrams, but only their topologies.
All topologies for two-point, three-point and the $\varphi^{Q}$-$Q\varphi$ correlation functions, as well as their cut graphs are generated by private Mathematica codes. The number of 1PI topologies for two-point and three-point functions, and irreducible correlation functions is listed in Table.(2).
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 1 - - - - 1 1 1
$L=2$ 5 2 - - - 7 2 3
$L=3$ 35 26 7 - - 68 9 17
$L=4$ 303 358 168 25 - 854 46 125
$L=5$ 3136 5300 3709 1103 128 13376 322 1170
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 1 - - - - 1 1 1
$L=2$ 7 4 - - - 11 2 5
$L=3$ 56 67 19 - - 142 9 35
$L=4$ 540 999 586 107 - 2232 46 303
$L=5$ 5905 15238 13846 5060 644 40693 322 3136
Table 2: (Top Table): the number of 1PI topologies of two-point, three-point correlation functions and the irreducible $\varphi^{Q}$-$Q\varphi$ correlation functions. (Bottom Table): the number of cut graphs, generated by the removal of one external leg for three-point function and (Q1)(Q-1) external legs for irreducible $\varphi^{Q}$-$Q\varphi$ correlation functions. These are the graphs to be computed by TwoPoint command for UV divergences.
The number grows rapidly as the number of loops increases. Cut graphs are generated through non-isomorphic computations of all possible external leg removals, and each of these cut graphs should be computed to determine their UV divergences. The number of cut graphs is listed in Table.(2). It can be seen that the number of cut graphs to be computed is about 3 to 4 times larger than the number of the original topologies.
The UV divergences are computed via TwoPoint command in HyperlogProcedures Maple package. The computational capability and time consumption of these evaluations are sensitive to both the loop order and the order of the dimensional regulator $\epsilon$. To apply the UV finiteness conditions for $L_{\tiny\mbox{max}}$-loop computation, an $L$-loop divergence in dimensional regularization should be computed up to the $\epsilon^{L_{\tiny\mbox{max}}-1-L}$ order. Increasing $L_{\tiny\mbox{max}}$ by one order requires UV divergence computations with an additional order in $\epsilon$ for all loop integrals, which significantly increases time consumption. For 5-loop computations, an $L$-loop divergence should be computed to the $\epsilon^{4-L}$ order, and the current version of HyperlogProcedures is capable of evaluating all required two-point integrals. Time consumption varies depending on the computer’s CPU and memory usage. On a normal desktop with a CPU base frequency 3.20GHz and 64GB memory, the approximate time required to evaluate all loop integrals is listed in Table.(3).
$Q=2$ $Q=3$ $Q=4$ $Q=5$ $Q=6$ Total 2-point 3-point
$L=1$ 0 - - - - 0 0 0
$L=2$ 0.02 h 0 - - - 0.02 h 0 0
$L=3$ 0.6 h 0.2 h 0 - - 0.8 h 0.03 h 0.15 h
$L=4$ 7 h 5.5 h 1.5 h 0 - 14 h 2.35 h 5.5 h
$L=5$ 16 d 24 d 10 d 1 d 0.2 h 51 d 5 d 8 d
Table 3: The approximate time consumption for 5-loop computations in a normal desktop. The symbols h and d are abbreviation of computer core hours and days. The working time less than 0.01 hour is simply denoted by 0.
Time consumption also increases rapidly as the loop order rises. However, for 5-loop computations, it remains within a reasonably acceptable range. In contrast, the symmetry factors and coupling factors for these topologies are generated almost instantaneously compared to the time required for evaluating UV divergences.
We have computed $Z_{\varphi}$, $Z_{\sigma}$, $Z_{h}$, $Z_{g}$ and all $\widetilde{Z}_{\varphi^{Q_{L}}},Q_{L}=2,\ldots,L+1$ up to 5-loops. From these results, we have derived the beta functions $\beta(g),\beta(h)$, the field anomalous dimensions $\gamma_{\varphi}$, $\gamma_{\sigma}$, and the scaling dimensions of the $\varphi^{Q}$} operator up to 5-loops. Results of lower than 5-loops have been compared with existing literature with agreements Fei:2014xta ; Gracey:2015tta ; Kompaniets:2021hwg .
The 5-loop correction to the anomalous dimension $\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}$ is presented in Appendix §A. The scaling dimension of the $\varphi^{Q}$ operator at the Wilson-Fisher fixed point can be computed as
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\left(\gamma_{\varphi^{Q}}+Q\gamma_{\varphi}\right)\Big{|}_{g=g^{\ast},h=h^{\ast}}~,$ (97)
where $g^{\ast},h^{\ast}$ are the fixed-point solutions of the beta functions. Our 5-loop fixed-point solutions, provided in Appendix §B, are in perfect agreement with the known results in the literature Kompaniets:2021hwg . By substituting these fixed-point solutions into the anomalous dimensions, we obtain the $\Delta_{Q}$ in the large $N$ limit. The leading, sub-leading and sub-sub-leading order contributions of the large $N$ expansion are as follows,
$\Delta_{Q\tiny\mbox{FP}}=\left(2-\frac{\epsilon}{2}\right)Q+\frac{Q}{N}\Delta_{Q\tiny\mbox{FP}}^{1}+\frac{Q}{N^{2}}\Delta_{Q\tiny\mbox{FP}}^{2}+\frac{Q}{N^{3}}\Delta_{Q\tiny\mbox{FP}}^{3}+\frac{Q}{N^{4}}\Delta_{Q\tiny\mbox{FP}}^{4}+\frac{Q}{N^{5}}\Delta_{Q\tiny\mbox{FP}}^{5}+\mathcal{O}(\frac{1}{N^{6}})~,$ (98)
where
$\Delta_{Q{\tiny\mbox{FP}}}^{1}$ $=$ $\bigg{[}4-3Q\bigg{]}\epsilon+\bigg{[}-\frac{8}{3}+\frac{7Q}{4}\bigg{]}\epsilon^{2}+\bigg{[}-\frac{7}{9}+\frac{11Q}{16}\bigg{]}\epsilon^{3}$ (99)
$+\bigg{[}\zeta_{3}-\frac{8}{27}+\left(\frac{19}{64}-\frac{3\zeta_{3}}{4}\right)Q\bigg{]}\epsilon^{4}+\bigg{[}\frac{\pi^{4}}{120}-\frac{43}{324}-\frac{2\zeta_{3}}{3}+\left(\frac{7\zeta_{3}}{16}-\frac{\pi^{4}}{160}+\frac{35}{256}\right)Q\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{2}$ $=$ $\bigg{[}176-132Q\bigg{]}\epsilon+\bigg{[}-\frac{1568}{3}+\frac{857Q}{2}-45Q^{2}\bigg{]}\epsilon^{2}$ (100)
$+\bigg{[}\frac{4105}{18}+72\zeta_{3}-\left(108\zeta_{3}+\frac{3743}{24}\right)Q+\left(36\zeta_{3}+\frac{93}{4}\right)Q^{2}\bigg{]}\epsilon^{3}$
$+\bigg{[}74\zeta_{3}+\frac{3\pi^{4}}{5}+\frac{19315}{216}+\left(21\zeta_{3}-\frac{9\pi^{4}}{10}-\frac{7465}{72}\right)Q+\left(\frac{3\pi^{4}}{10}-51\zeta_{3}+\frac{167}{8}\right)Q^{2}\bigg{]}\epsilon^{4}$
$+\bigg{[}\frac{12641}{324}-\frac{2281\zeta_{3}}{6}+\frac{37\pi^{4}}{60}+36\zeta_{5}+\left(\frac{577\zeta_{3}}{2}-54\zeta_{5}+\frac{7\pi^{4}}{40}-\frac{44459}{864}\right)Q$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(18\zeta_{5}-\frac{65\zeta_{3}}{4}+\frac{339}{32}-\frac{17\pi^{4}}{40}\right)Q^{2}\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{3}$ $=$ $\bigg{[}7744-5808Q\bigg{]}\epsilon+\bigg{[}-66872+59520Q-9000Q^{2}\bigg{]}\epsilon^{2}$ (101)
$+\bigg{[}\frac{193141}{2}-4212\zeta_{3}-\bigg{(}3996\zeta_{3}+88575\bigg{)}Q+\bigg{(}4536\zeta_{3}+20538\bigg{)}Q^{2}-1350Q^{3}\bigg{]}\epsilon^{3}$
$+\bigg{[}61920\zeta_{3}-14040\zeta_{5}-\frac{351\pi^{4}}{10}-\frac{107537}{216}+\bigg{(}14580\zeta_{5}-22146\zeta_{3}-\frac{63913}{6}-\frac{333\pi^{4}}{10}\bigg{)}Q$
$~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{2101}{2}-18558\zeta_{3}-1080\zeta_{5}+\frac{189\pi^{4}}{5}\bigg{)}Q^{2}+\bigg{(}\frac{1107}{2}+1728\zeta_{3}+540\zeta_{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{4}$
$+\bigg{[}516\pi^{4}-540\zeta_{3}^{2}-166574\zeta_{3}-\frac{35397\zeta_{5}}{2}-\frac{130\pi^{6}}{7}-\frac{14477777}{864}$
$~~~~~~~~+\bigg{(}\frac{9276203}{432}-2106\zeta_{3}^{2}+\frac{245019\zeta_{3}}{2}-\frac{981\zeta_{5}}{2}+\frac{135\pi^{6}}{7}-\frac{3691\pi^{4}}{20}\bigg{)}Q$
$~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}3348\zeta_{3}^{2}+\frac{549\zeta_{3}}{2}+3663\zeta_{5}-\frac{10\pi^{6}}{7}-\frac{3093\pi^{4}}{20}-\frac{190189}{24}\bigg{)}Q^{2}$
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\bigg{(}\frac{7551}{8}-702\zeta_{3}^{2}-2322\zeta_{3}-1215\zeta_{5}+\frac{5\pi^{6}}{7}+\frac{72\pi^{4}}{5}\bigg{)}Q^{3}\bigg{]}\epsilon^{5}~,$
and expressions $\Delta_{Q{\tiny\mbox{FP}}}^{4},\Delta_{Q{\tiny\mbox{FP}}}^{5}$ are presented in Appendix §C. As a confirmation of these results, the 3-loop contribution up to $\epsilon^{3}$ order is consistent with literature findings Vasiliev:1981yc ; Vasiliev:1981dg ; Vasiliev:1982dc ; Jack:2021ziq  [^7] The value $a_{t}^{(3)}$ in paper Jack:2021ziq that computes $\Delta_{Q}^{(3)}$ is incorrect. After using the correct value $a_{t}^{(3)}=\frac{g^{6}}{3}+g^{5}h$, the $\Delta_{Q}^{(3)}$ and the $1/N^{3}$ term in their paper are consistent with the 3-loop order part of our results.. The corresponding terms $\frac{Q^{i+1}\epsilon^{i}}{N^{i}},\frac{Q^{i+1}\epsilon^{i+1}}{N^{i}},i=1,\ldots,5$ are also consistent with those reported in the semi-classical literature Antipin:2021jiw . The 4-loop contribution up to $1/N^{2}$ order has been previously provided in Huang:2024hsn , while the 5-loop contributions up to $1/N^{5}$ order presented here are new. Our results have been provided in ancillary files to the arXiv submission of this article.
5 Conclusion
In this paper, we apply the OPE method of renormalization proposed in Huang:2024hsn to evaluate the 5-loop scaling dimensions of the $\phi^{Q}$ operator in the six-dimensional cubic scalar theory. The result is derived from over 40 000 two-point propagator-type integrals computed using the graphical function method, combined with non-trivial summation among these integrals. We have successfully obtained the 5-loop correction to the anomalous dimension of the $\phi^{Q}$ operator, as well as the large $N$ expansion of the scaling dimension at the Wilson-Fisher fixed point up to the $1/N^{5}$ order. It stands for a new record for the scaling dimensions of $\phi^{Q}$ operator in cubic scalar theory. The result shows consistency with lower-loop perturbative calculations reported in other literature and aligns with the corresponding terms derived via semi-classical methods. Furthermore, the non-trivial cancellation of $\epsilon$ poles in the final result provides a strong consistency check validating its correctness. This work further demonstrates the efficiency and power of the OPE method in renormalization computations.
Efforts to apply the OPE method to address the challenges posed by more complex problems are currently underway. A promising project aims to surpass the 6-loop record in quartic scalar theory by computing the 7-loop scaling dimensions of the $\phi^{Q}$ operator. Such research would not only advance the precision of renormalization computations but also drive the evolution of modern techniques for multi-loop integral evaluations. Also as previously noted, the OPE method is highly general and extendable beyond scalar theories. It is interesting to investigate the renormalization computations by OPE method in other situations, such as the Gross-Neveu-Yukawa theory with fermions, or scalar-QED theory with gauge fields, etc. In these theories, Feynman integrals are no longer purely scalar integrals, but involve non-trivial numerators. While further developments may enable the computation of such non-trivial master integrals via graphical function method adapted to numerator structures, the most viable current approach relies on the computational algorithms Smirnov:2008iw ; Maierhofer:2017gsa ; Wu:2023upw ; Guan:2024byi based on the IBP method Tkachov:1981wb ; Chetyrkin:1981qh ; Laporta:2000dsw . Another potential challenge would be applying the OPE method to the anomalous dimensions of more intricate operators, such as those with higher derivatives, Lorentz indices, or in scenarios with operator mixing Derkachov:1997ch ; Cao:2021cdt . While the application of the OPE method remains straightforward in these cases Huang:2024hsn , practical computations demand more sophisticated techniques. It is interesting to explore the diverse applications of the OPE method in renormalization, and further enhance its capabilities.
Acknowledgments
We would like to thank Oliver Schnetz for stimulating conversations on the graphical function method and the usage of HyperlogProcedures package. This work is supported by the Science Challenge Project (No. TZ2025012), and NSAF No.U2330401.
Appendix A Five-loop corrections to the scaling dimensions
In this section, we collect our results of the scaling dimension $\Delta_{Q}$ in six-dimensional cubic scalar theory. Note that the 4-loop result has been presented in Huang:2024hsn . The 5-loop correction of $\Delta_{Q}$ is given by
$\Delta_{Q}^{{\tiny\mbox{5-loop}}}=Q\sum_{i=2}^{10}g^{i}h^{10-i}\delta^{5}_{i}~,$ (102)
where,
$\delta^{5}_{2}$ $=$ $-\frac{3336089}{13436928}+\frac{25882183Q}{161243136}+\left(-\frac{685}{7776}+\frac{15341Q}{248832}\right)\zeta_{3}+\left(-\frac{1177}{1866240}+\frac{23Q}{55296}\right)\pi^{4}$ (103)
$~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\left(\frac{311}{1296}-\frac{271Q}{1728}\right)\zeta_{5}+\left(\frac{5}{40824}-\frac{5Q}{54432}\right)\pi^{6}+\left(-\frac{1}{108}+\frac{Q}{144}\right)\zeta_{3}^{~2}~,$
$\delta^{5}_{3}$ $=$ $-\frac{133429553}{161243136}+\frac{26603591Q}{13436928}-\frac{228283Q^{2}}{279936}+\left(\frac{231613}{62208}-\frac{28829Q}{6912}+\frac{13091Q^{2}}{15552}\right)\zeta_{3}+\left(-\frac{241}{46080}\right$. (104)
$\left.+\frac{121Q}{11520}-\frac{2599Q^{2}}{622080}\right)\pi^{4}+\left(-\frac{29}{16}+\frac{421Q}{576}+\frac{719Q^{2}}{1728}\right)\zeta_{5}+\left(-\frac{5}{3402}+\frac{5Q}{4536}\right)\pi^{6}+\left(\frac{1}{9}-\frac{Q}{12}\right)\zeta_{3}^{~2}~,$
$\delta^{5}_{4}$ $=$ $\frac{4152458077}{241864704}+\frac{1256311N}{20155392}-\frac{1679472427Q}{80621568}-\frac{2125781NQ}{53747712}+\frac{24768415Q^{2}}{4478976}-\frac{20291Q^{3}}{41472}$ (105)
$+\left(\frac{104597}{3456}-\frac{65N}{2592}-\frac{2478701Q}{62208}+\frac{3919NQ}{248832}+\frac{169543Q^{2}}{15552}-\frac{131Q^{3}}{144}\right)\zeta_{3}+\left(-\frac{31489}{1866240}\right$.
$\left.-\frac{11N}{622080}+\frac{307Q}{13824}+\frac{11NQ}{497664}-\frac{2023Q^{2}}{311040}+\frac{Q^{3}}{1080}\right)\pi^{4}+\left(-\frac{33983}{2592}+\frac{N}{324}+\frac{6473Q}{288}-\frac{NQ}{432}\right$.
$\left.-\frac{3041Q^{2}}{288}+\frac{85Q^{3}}{72}\right)\zeta_{5}+\left(\frac{205}{20412}-\frac{25Q}{1944}+\frac{25Q^{2}}{6804}-\frac{5Q^{3}}{13608}\right)\pi^{6}+\left(\frac{1195}{216}-\frac{323Q}{72}-\frac{37Q^{2}}{36}\right$.
$\left.+\frac{4Q^{3}}{9}\right)\zeta_{3}^{~2}+\left(-\frac{5467}{96}+\frac{6055Q}{96}-\frac{147Q^{2}}{16}\right)\zeta_{7}~,$
$\delta^{5}_{5}$ $=$ $-\frac{705928679}{26873856}+\frac{25722139N}{161243136}+\frac{519118363Q}{13436928}-\frac{3257NQ}{7776}-\frac{38531567Q^{2}}{2239488}+\frac{129851NQ^{2}}{746496}$ (106)
$+\frac{10871Q^{3}}{3072}-\frac{641Q^{4}}{2304}+\left(-\frac{1604333}{31104}-\frac{1145N}{1536}+\frac{581737Q}{6912}+\frac{23861NQ}{20736}-\frac{1196249Q^{2}}{31104}\right$.
$\left.-\frac{3719NQ^{2}}{10368}+\frac{7727Q^{3}}{1152}-\frac{11Q^{4}}{24}\right)\zeta_{3}+\left(\frac{4411}{155520}+\frac{607N}{124416}-\frac{11137Q}{207360}-\frac{17NQ}{2592}+\frac{1997Q^{2}}{62208}\right$.
$\left.+\frac{43NQ^{2}}{23040}-\frac{161Q^{3}}{34560}\right)\pi^{4}+\left(\frac{12157}{576}+\frac{43N}{216}-\frac{15383Q}{576}-\frac{49NQ}{192}+\frac{9767Q^{2}}{1728}+\frac{107NQ^{2}}{1728}\right$.
$\left.+\frac{5Q^{3}}{9}-\frac{Q^{4}}{3}\right)\zeta_{5}+\left(\frac{305}{54432}+\frac{25Q}{36288}-\frac{95Q^{2}}{15552}+\frac{5Q^{3}}{3402}\right)\pi^{6}+\left(-\frac{223}{144}-\frac{371Q}{96}+\frac{1723Q^{2}}{288}\right$.
$\left.-\frac{10Q^{3}}{9}\right)\zeta_{3}^{~2}+\left(\frac{917}{16}-\frac{581Q}{6}+\frac{2401Q^{2}}{48}-\frac{35Q^{3}}{3}+\frac{7Q^{4}}{6}\right)\zeta_{7}~,$
$\delta^{5}_{6}$ $=$ $\frac{101332127}{1889568}-\frac{377140433N}{80621568}+\frac{20375N^{2}}{13436928}-\frac{151156525Q}{1679616}+\frac{923790881NQ}{161243136}-\frac{84245N^{2}Q}{80621568}+\frac{3532885Q^{2}}{69984}$ (107)
$-\frac{3589273NQ^{2}}{2239488}-\frac{14759Q^{3}}{1024}+\frac{9473NQ^{3}}{82944}+\frac{545Q^{4}}{256}-\frac{7Q^{5}}{48}+\left(\frac{1035907}{7776}+\frac{4147N}{62208}+\frac{41N^{2}}{46656}-\frac{6811451Q}{31104}\right$.
$\left.+\frac{44599NQ}{62208}-\frac{197N^{2}Q}{248832}+\frac{1258969Q^{2}}{10368}-\frac{36589NQ^{2}}{62208}-\frac{19235Q^{3}}{576}+\frac{17NQ^{3}}{144}+\frac{65Q^{4}}{24}\right)\zeta_{3}$
$+\left(\frac{50047}{466560}-\frac{719N}{311040}-\frac{17N^{2}}{1866240}-\frac{35393Q}{311040}+\frac{7NQ}{3072}+\frac{17N^{2}Q}{2488320}+\frac{349Q^{2}}{20736}-\frac{103NQ^{2}}{622080}+\frac{73Q^{3}}{17280}-\frac{NQ^{3}}{2160}\right)\pi^{4}$
$+\left(\frac{26623}{144}+\frac{7325N}{648}-\frac{296567Q}{864}-\frac{25771NQ}{1728}+\frac{186683Q^{2}}{864}+\frac{851NQ^{2}}{192}-\frac{1435Q^{3}}{24}-\frac{65NQ^{3}}{144}+\frac{35Q^{4}}{6}\right)\zeta_{5}$
$+\left(\frac{415}{163296}-\frac{185N}{40824}+\frac{5Q}{2268}+\frac{55NQ}{9072}-\frac{295Q^{2}}{54432}-\frac{25NQ^{2}}{13608}+\frac{25Q^{3}}{13608}+\frac{5NQ^{3}}{27216}\right)\pi^{6}$
$+\left(-\frac{15473}{432}-\frac{367N}{216}+71Q+\frac{53NQ}{24}-\frac{5899Q^{2}}{144}-\frac{11NQ^{2}}{18}+\frac{223Q^{3}}{36}+\frac{NQ^{3}}{36}\right)\zeta_{3}^{~2}$
$+\left(-\frac{8827}{24}+\frac{19817Q}{32}-\frac{17129Q^{2}}{48}+\frac{413Q^{3}}{4}-\frac{35Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{7}$ $=$ $-\frac{1727695009}{13436928}+\frac{1219348333N}{80621568}-\frac{2802343N^{2}}{53747712}+\frac{1827570041Q}{6718464}-\frac{38406623NQ}{2239488}+\frac{42845N^{2}Q}{2239488}$ (108)
$-\frac{360106991Q^{2}}{2239488}+\frac{22560479NQ^{2}}{4478976}+\frac{35173N^{2}Q^{2}}{4478976}+\frac{1125739Q^{3}}{41472}-\frac{20573NQ^{3}}{20736}+\frac{3223Q^{4}}{4608}+\frac{133NQ^{4}}{2304}-\frac{7Q^{5}}{16}$
$+\left(-\frac{48245}{3888}+\frac{1642177N}{31104}+\frac{3011N^{2}}{62208}+\frac{128345Q}{972}-\frac{4686905NQ}{62208}-\frac{331N^{2}Q}{5184}-\frac{1930387Q^{2}}{10368}+\frac{406627NQ^{2}}{15552}\right$.
$\left.+\frac{1165N^{2}Q^{2}}{62208}+\frac{14215Q^{3}}{192}-\frac{1451NQ^{3}}{576}-\frac{141Q^{4}}{16}\right)\zeta_{3}+\left(-\frac{177979}{311040}-\frac{71161N}{622080}-\frac{95N^{2}}{248832}+\frac{162677Q}{155520}\right$.
$\left.+\frac{103393NQ}{622080}+\frac{17N^{2}Q}{23040}-\frac{19751Q^{2}}{34560}-\frac{3713NQ^{2}}{62208}-\frac{91N^{2}Q^{2}}{311040}+\frac{1781Q^{3}}{17280}+\frac{89NQ^{3}}{17280}\right)\pi^{4}+\left(-\frac{2030531}{1728}\right$.
$\left.-\frac{47359N}{576}-\frac{N^{2}}{144}+\frac{1301549Q}{576}+\frac{203777NQ}{1728}+\frac{N^{2}Q}{96}-\frac{824411Q^{2}}{576}-\frac{26117NQ^{2}}{576}-\frac{N^{2}Q^{2}}{288}+\frac{6655Q^{3}}{18}\right$.
$\left.+\frac{445NQ^{3}}{72}-\frac{110Q^{4}}{3}\right)\zeta_{5}+\left(\frac{3725}{27216}+\frac{55N}{3402}-\frac{8185Q}{36288}-\frac{3085NQ}{108864}+\frac{11695Q^{2}}{108864}+\frac{1405NQ^{2}}{108864}-\frac{205Q^{3}}{13608}\right$.
$\left.-\frac{25NQ^{3}}{13608}\right)\pi^{6}+\left(\frac{9959}{72}+\frac{629N}{72}-\frac{22685Q}{96}-\frac{3493NQ}{288}+\frac{34195Q^{2}}{288}+\frac{1273NQ^{2}}{288}-\frac{685Q^{3}}{36}-\frac{11NQ^{3}}{18}\right)\zeta_{3}^{~2}$
$+\left(\frac{12887}{12}-\frac{147N}{16}-\frac{18291Q}{8}+\frac{147NQ}{16}+\frac{9793Q^{2}}{6}-\frac{3731Q^{3}}{8}+\frac{140Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{8}$ $=$ $\frac{26586982301}{60466176}-\frac{779370073N}{60466176}+\frac{26353453N^{2}}{80621568}-\frac{3827N^{3}}{20155392}-\frac{8024021965Q}{10077696}+\frac{79115735NQ}{5038848}+\frac{1162525N^{2}Q}{20155392}$ (109)
$+\frac{2029N^{3}Q}{17915904}+\frac{514446007Q^{2}}{1119744}-\frac{4617385NQ^{2}}{559872}-\frac{25159N^{2}Q^{2}}{165888}-\frac{2234915Q^{3}}{20736}+\frac{237013NQ^{3}}{82944}+\frac{127N^{2}Q^{3}}{27648}$
$+\frac{15847Q^{4}}{1536}-\frac{163NQ^{4}}{768}-\frac{3Q^{5}}{4}+\left(-\frac{8364127}{15552}-\frac{929849N}{23328}-\frac{18653N^{2}}{31104}-\frac{43N^{3}}{31104}+\frac{7945621Q}{7776}+\frac{2483159NQ}{31104}\right$.
$\left.+\frac{33251N^{2}Q}{62208}+\frac{245N^{3}Q}{248832}-\frac{8598169Q^{2}}{15552}-\frac{163897NQ^{2}}{3456}-\frac{1465N^{2}Q^{2}}{20736}+\frac{9743Q^{3}}{108}+\frac{899NQ^{3}}{144}-\frac{149Q^{4}}{24}\right$.
$\left.+\frac{3NQ^{4}}{8}\right)\zeta_{3}+\left(\frac{18179}{15552}+\frac{89461N}{311040}+\frac{2899N^{2}}{622080}+\frac{N^{3}}{622080}-\frac{314647Q}{155520}-\frac{145487NQ}{311040}-\frac{4417N^{2}Q}{622080}-\frac{N^{3}Q}{829440}\right$.
$\left.+\frac{31291Q^{2}}{31104}+\frac{7607NQ^{2}}{34560}+\frac{481N^{2}Q^{2}}{207360}-\frac{151Q^{3}}{1152}-\frac{29NQ^{3}}{864}\right)\pi^{4}+\left(\frac{6140413}{2592}+\frac{784319N}{5184}+\frac{43N^{2}}{864}-\frac{4206853Q}{864}\right$.
$\left.-\frac{406439NQ}{1728}-\frac{7N^{2}Q}{64}+\frac{2890897Q^{2}}{864}+\frac{7891NQ^{2}}{72}+\frac{103N^{2}Q^{2}}{1728}-\frac{44575Q^{3}}{48}-\frac{1735NQ^{3}}{96}+\frac{260Q^{4}}{3}\right)\zeta_{5}$
$+\left(-\frac{9425}{40824}-\frac{7015N}{163296}+\frac{7565Q}{18144}+\frac{4255NQ}{54432}-\frac{3995Q^{2}}{18144}-\frac{325NQ^{2}}{7776}+\frac{965Q^{3}}{27216}+\frac{365NQ^{3}}{54432}\right)\pi^{6}$
$+\left(-\frac{5089}{216}-\frac{5359N}{432}+\frac{2083Q}{48}+\frac{4207NQ}{144}-\frac{1003Q^{2}}{48}-\frac{2935NQ^{2}}{144}+\frac{155Q^{3}}{72}+\frac{581NQ^{3}}{144}\right)\zeta_{3}^{~2}$
$+\left(-\frac{200585}{96}-\frac{1631N}{24}+\frac{415247Q}{96}+\frac{371NQ}{6}-\frac{148813Q^{2}}{48}+\frac{11123Q^{3}}{12}-\frac{280Q^{4}}{3}\right)\zeta_{7}~,$
$\delta^{5}_{9}$ $=$ $-\frac{4988268911}{10077696}+\frac{158410849N}{4478976}-\frac{1951205N^{2}}{20155392}+\frac{162107N^{3}}{161243136}+\frac{2881576087Q}{3359232}-\frac{213374875NQ}{6718464}-\frac{720205N^{2}Q}{746496}$ (110)
$-\frac{1213N^{3}Q}{2239488}-\frac{237912325Q^{2}}{559872}+\frac{2904071NQ^{2}}{746496}+\frac{2387851N^{2}Q^{2}}{4478976}+\frac{11N^{3}Q^{2}}{1492992}+\frac{146737Q^{3}}{2592}+\frac{2015NQ^{3}}{13824}-\frac{209N^{2}Q^{3}}{82944}$
$+\frac{673Q^{4}}{144}-\frac{1915NQ^{4}}{4608}-\frac{5Q^{5}}{6}+\left(\frac{11111}{32}+\frac{992585N}{15552}+\frac{47023N^{2}}{31104}+\frac{1397N^{3}}{124416}-\frac{14576275Q}{15552}-\frac{6227NQ}{54}-\frac{5197N^{2}Q}{7776}\right$.
$\left.-\frac{23N^{3}Q}{3456}+\frac{2176717Q^{2}}{2592}+\frac{674003NQ^{2}}{10368}-\frac{659N^{2}Q^{2}}{10368}-\frac{7N^{3}Q^{2}}{6912}-\frac{39937Q^{3}}{144}-\frac{6193NQ^{3}}{576}-\frac{35N^{2}Q^{3}}{384}+\frac{215Q^{4}}{8}\right$.
$\left.-\frac{5NQ^{4}}{24}\right)\zeta_{3}+\left(\frac{1703}{5184}-\frac{30931N}{62208}-\frac{203N^{2}}{69120}-\frac{N^{3}}{51840}-\frac{12353Q}{31104}+\frac{4771NQ}{5760}+\frac{233N^{2}Q}{31104}+\frac{N^{3}Q}{69120}+\frac{1427Q^{2}}{25920}\right$.
$\left.-\frac{511NQ^{2}}{1280}-\frac{41N^{2}Q^{2}}{10368}+\frac{7Q^{3}}{240}+\frac{73NQ^{3}}{1152}+\frac{N^{2}Q^{3}}{2304}\right)\pi^{4}+\left(-\frac{1144559}{432}-\frac{321965N}{1728}-\frac{67N^{2}}{96}+\frac{599057Q}{108}\right$.
$\left.+\frac{534791NQ}{1728}+\frac{847N^{2}Q}{864}-\frac{1665455Q^{2}}{432}-\frac{93157NQ^{2}}{576}-\frac{37N^{2}Q^{2}}{72}+\frac{76435Q^{3}}{72}+\frac{235NQ^{3}}{9}-105Q^{4}+\frac{5NQ^{4}}{6}\right)\zeta_{5}$
$+\left(-\frac{155}{6804}+\frac{2735N}{27216}-\frac{25N^{2}}{54432}+\frac{55Q}{1512}-\frac{18325NQ}{108864}+\frac{25N^{2}Q}{54432}-\frac{5Q^{2}}{432}+\frac{2995NQ^{2}}{36288}+\frac{5Q^{3}}{2268}-\frac{5NQ^{3}}{378}\right)\pi^{6}$
$+\left(-\frac{3037}{9}+\frac{2369N}{72}-\frac{19N^{2}}{144}+\frac{6817Q}{12}-\frac{16729NQ}{288}+\frac{19N^{2}Q}{144}-\frac{6445Q^{2}}{24}+\frac{2951NQ^{2}}{96}+\frac{112Q^{3}}{3}-5NQ^{3}\right)\zeta_{3}^{~2}$
$+\left(\frac{78659}{24}-\frac{203N}{12}-\frac{299551Q}{48}+\frac{259NQ}{16}+\frac{181027Q^{2}}{48}+\frac{35NQ^{2}}{48}-\frac{21077Q^{3}}{24}+70Q^{4}\right)\zeta_{7}~,$
$\delta^{5}_{10}$ $=$ $\frac{3873526685}{15116544}-\frac{1261849117N}{60466176}-\frac{1111975N^{2}}{30233088}+\frac{2132959N^{3}}{241864704}-\frac{43N^{4}}{2519424}-\frac{4120586761Q}{10077696}+\frac{217226153NQ}{6718464}$ (111)
$+\frac{24155359N^{2}Q}{26873856}-\frac{1574323N^{3}Q}{161243136}+\frac{35N^{4}Q}{1990656}+\frac{12775867Q^{2}}{69984}-\frac{16632703NQ^{2}}{1119744}-\frac{239203N^{2}Q^{2}}{559872}+\frac{41N^{3}Q^{2}}{31104}$
$-\frac{3373Q^{3}}{144}+\frac{3283NQ^{3}}{1152}-\frac{133N^{2}Q^{3}}{4608}+\frac{7Q^{4}}{64}-\frac{469NQ^{4}}{1536}-\frac{Q^{5}}{2}+\left(\frac{2181835}{11664}-\frac{33815N}{1458}-\frac{246587N^{2}}{46656}-\frac{3619N^{3}}{186624}\right$.
$\left.-\frac{N^{4}}{11664}-\frac{2534795Q}{15552}+\frac{502549NQ}{5184}+\frac{119575N^{2}Q}{20736}+\frac{37N^{3}Q}{4608}+\frac{7N^{4}Q}{124416}-\frac{9701Q^{2}}{81}-\frac{390451NQ^{2}}{5184}-\frac{11453N^{2}Q^{2}}{7776}\right$.
$\left.+\frac{31N^{3}Q^{2}}{7776}+\frac{50953Q^{3}}{432}+\frac{4409NQ^{3}}{432}+\frac{143N^{2}Q^{3}}{576}-\frac{47Q^{4}}{2}-\frac{19NQ^{4}}{48}\right)\zeta_{3}+\left(-\frac{10277}{23328}+\frac{17249N}{58320}+\frac{667N^{2}}{933120}\right$.
$\left.-\frac{23N^{3}}{233280}+\frac{N^{4}}{933120}+\frac{27649Q}{31104}-\frac{153709NQ}{311040}-\frac{1493N^{2}Q}{311040}+\frac{97N^{3}Q}{622080}-\frac{N^{4}Q}{1244160}-\frac{199Q^{2}}{360}+\frac{13021NQ^{2}}{51840}\right$.
$\left.+\frac{137N^{2}Q^{2}}{38880}-\frac{17N^{3}Q^{2}}{311040}+\frac{77Q^{3}}{720}-\frac{137NQ^{3}}{3456}-\frac{13N^{2}Q^{3}}{17280}\right)\pi^{4}+\left(\frac{185473}{144}+\frac{88807N}{864}+\frac{28717N^{2}}{5184}+\frac{N^{3}}{216}\right$.
$\left.-\frac{378875Q}{144}-\frac{345655NQ}{1728}-\frac{3679N^{2}Q}{576}-\frac{N^{3}Q}{144}+\frac{387313Q^{2}}{216}+\frac{99619NQ^{2}}{864}+\frac{475N^{2}Q^{2}}{288}+\frac{N^{3}Q^{2}}{432}-\frac{3915Q^{3}}{8}\right$.
$\left.-\frac{2825NQ^{3}}{144}-\frac{5N^{2}Q^{3}}{32}+\frac{125Q^{4}}{3}-\frac{NQ^{4}}{3}\right)\zeta_{5}+\left(\frac{445}{3402}-\frac{145N}{2268}-\frac{3155Q}{13608}+\frac{6025NQ}{54432}+\frac{10Q^{2}}{81}-\frac{3055NQ^{2}}{54432}\right$.
$\left.-\frac{5N^{2}Q^{2}}{27216}-\frac{65Q^{3}}{3402}+\frac{25NQ^{3}}{3024}+\frac{5N^{2}Q^{3}}{54432}\right)\pi^{6}+\left(\frac{2342}{9}-\frac{209N}{24}+\frac{23N^{2}}{36}-\frac{4145Q}{9}+\frac{3271NQ}{144}-N^{2}Q+240Q^{2}\right$.
$\left.-\frac{2287NQ^{2}}{144}+\frac{31N^{2}Q^{2}}{72}-\frac{344Q^{3}}{9}+\frac{49NQ^{3}}{24}-\frac{13N^{2}Q^{3}}{144}\right)\zeta_{3}^{~2}+\left(-\frac{216181}{96}-\frac{567N}{32}+\frac{131411Q}{32}-\frac{105NQ}{4}\right$.
$\left.-2296Q^{2}+\frac{301NQ^{2}}{8}+\frac{3563Q^{3}}{8}-\frac{119NQ^{3}}{48}-\frac{49Q^{4}}{3}+\frac{7NQ^{4}}{6}\right)\zeta_{7}~.$
Appendix B Solutions at the Wilson-Fisher fixed point
The beta functions can be solved order by order in $\epsilon$, which leads to the solutions at fixed point as,
$g^{\ast}$ $=$ $\sqrt{\frac{6\epsilon}{N}}\bigg{\{}1+\frac{22}{N}+\frac{726}{N^{2}}-\frac{326180}{N^{3}}-\frac{349658330}{N^{4}}+\bigg{[}-\frac{155}{6N}-\frac{1705}{N^{2}}+\frac{912545}{N^{3}}+\frac{3590574890}{3N^{4}}\bigg{]}\epsilon+\bigg{[}\frac{1777}{144N}$ (112)
$+\frac{1}{N^{2}}\bigg{(}\frac{29093}{36}-1170\zeta_{3}\bigg{)}+\frac{1}{N^{3}}\bigg{(}-\frac{106755739}{72}-228060\zeta_{3}\bigg{)}+\frac{92543220\zeta_{3}-1920046163}{N^{4}}\bigg{]}\epsilon^{2}+\bigg{[}-\frac{217}{2592N}$
$+\frac{1}{N^{2}}\bigg{(}\frac{709151}{864}+3126\zeta_{3}-\frac{13\pi^{4}}{2}+360\zeta_{5}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{779869165}{432}+2229915\zeta_{3}-1267\pi^{4}-2438640\zeta_{5}\bigg{)}$
$+\frac{1}{N^{4}}\bigg{(}\frac{2437481884309}{1296}+278186790\zeta_{3}+864409\pi^{4}-1469512080\zeta_{5}\bigg{)}\bigg{]}\epsilon^{3}+\bigg{[}\frac{1}{N}\bigg{(}\frac{6973}{41472}-\frac{155\zeta_{3}}{96}\bigg{)}$
$+\frac{1}{N^{2}}\bigg{(}-\frac{49050023}{124416}-\frac{40871\zeta_{3}}{24}+\frac{1563\pi^{4}}{80}-4320\zeta_{5}+\frac{5\pi^{6}}{14}+81\zeta_{3}^{2}\bigg{)}+\frac{1}{N^{3}}\bigg{(}-\frac{17059272503}{15552}-\frac{77297317\zeta_{3}}{48}$
$+\frac{667967\pi^{4}}{48}+\frac{23014935\zeta_{5}}{4}-\frac{16935\pi^{6}}{7}-366930\zeta_{3}^{2}-\frac{2941029\zeta_{7}}{2}\bigg{)}+\frac{1}{N^{4}}\bigg{(}-\frac{68647291122307}{62208}+\frac{40450011887\zeta_{3}}{12}$
$-\frac{1345559\pi^{4}}{24}+\frac{13461052935\zeta_{5}}{2}-\frac{9381745\pi^{6}}{7}-37063008\zeta_{3}^{2}-6633588717\zeta_{7}\bigg{)}\bigg{]}\epsilon^{4}\bigg{\}}~,$
and
$h^{\ast}$ $=$ $6\sqrt{\frac{6\epsilon}{N}}\bigg{\{}1+\frac{162}{N}+\frac{68766}{N^{2}}+\frac{41224420}{N^{3}}+\frac{28762554870}{N^{4}}+\bigg{[}-\frac{215}{2N}-\frac{86335}{N^{2}}-\frac{75722265}{N^{3}}-\frac{69633402510}{N^{4}}\bigg{]}\epsilon+\bigg{[}\frac{927}{16N}$ (113)
$+\frac{1}{N^{2}}\bigg{(}\frac{270911}{6}-26190\zeta_{3}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{1156981601}{24}-26334180\zeta_{3}\bigg{)}+\frac{1}{N^{4}}\bigg{(}\frac{116908872997}{2}-23638808340\zeta_{3}\bigg{)}\bigg{]}\epsilon^{2}$
$+\bigg{[}-\frac{3461}{864N}+\frac{1}{N^{2}}\bigg{(}\frac{25209239}{864}+115860\zeta_{3}-\frac{291\pi^{4}}{2}-70200\zeta_{5}\bigg{)}+\frac{1}{N^{3}}\bigg{(}\frac{18166643735}{432}+198108185\zeta_{3}-175491\pi^{4}$
$-169538400\zeta_{5}\bigg{)}+\frac{1}{N^{4}}\bigg{(}\frac{5465613593485}{144}+235067598534\zeta_{3}-172334673\pi^{4}-211183698960\zeta_{5}\bigg{)}\bigg{]}\epsilon^{3}$
$+\bigg{[}\frac{1}{N}\bigg{(}\frac{5945}{13824}-\frac{215\zeta_{3}}{32}\bigg{)}+\frac{1}{N^{2}}\bigg{(}-\frac{1204206197}{41472}-\frac{973091\zeta_{3}}{8}+\frac{5793\pi^{4}}{8}+104895\zeta_{5}-\frac{975\pi^{6}}{14}-15795\zeta_{3}^{2}\bigg{)}+$
$\frac{1}{N^{3}}\bigg{(}-\frac{35896723781}{648}-\frac{1912729075\zeta_{3}}{16}+\frac{10821883\pi^{4}}{8}+\frac{2017087425\zeta_{5}}{4}-\frac{1245950\pi^{6}}{7}-23961852\zeta_{3}^{2}-\frac{637290423\zeta_{7}}{2}\bigg{)}$
$+\frac{1}{N^{4}}\bigg{(}-\frac{179864809668917}{2304}-\frac{552249422829\zeta_{3}}{4}+\frac{34006612103\pi^{4}}{20}+\frac{1850885090325\zeta_{5}}{2}-\frac{1654305365\pi^{6}}{7}$
$-15580792368\zeta_{3}^{2}-725681165454\zeta_{7}\bigg{)}\bigg{]}\epsilon^{4}\bigg{\}}~.$
Appendix C Large $N$ expansion at the Wilson-Fisher fixed point
Here we present the results of the $1/N^{4}$ and $1/N^{5}$ orders in the large $N$ expansion of $\gamma_{Q}$ at Wilson-Fisher fixed point. We have,
$\Delta_{Q{\tiny\mbox{FP}}}^{4}$ $=$ $\bigg{[}-2481664+1861248Q\bigg{]}\epsilon+\bigg{[}-\frac{21069760}{3}+8828264Q-3043440Q^{2}\bigg{]}\epsilon^{2}+\bigg{[}\frac{259232542}{9}-861840\zeta_{3}+\bigg{(}-30411798$ (114)
$-1662768\zeta_{3}\bigg{)}Q+\bigg{(}8984340+1478304\zeta_{3}\bigg{)}Q^{2}-495720Q^{3}\bigg{]}\epsilon^{3}+\bigg{[}-\frac{25851511}{4}+44545394\zeta_{3}-7182\pi^{4}-40335840\zeta_{5}$
$+\bigg{(}\frac{124128583}{24}-31054098\zeta_{3}-\frac{69282\pi^{4}}{5}+38233080\zeta_{5}\bigg{)}Q+\bigg{(}-4410989-2172420\zeta_{3}+\frac{61596\pi^{4}}{5}-4110480\zeta_{5}\bigg{)}Q^{2}$
$+\bigg{(}1024488+388854\zeta_{3}+186840\zeta_{5}\bigg{)}Q^{3}-\frac{213597Q^{4}}{4}\bigg{]}\epsilon^{4}+\bigg{[}-\frac{29834168837}{1296}-\frac{498959309\zeta_{3}}{6}+\frac{22272697\pi^{4}}{60}$
$+110252862\zeta_{5}-9004500\zeta_{3}^{2}-\frac{373480\pi^{6}}{7}-47146617\zeta_{7}+\bigg{(}\frac{23354482969}{864}+48023731\zeta_{3}-\frac{5175683\pi^{4}}{20}-89443629\zeta_{5}$
$+\frac{354010\pi^{6}}{7}+10983168\zeta_{3}^{2}+42115059\zeta_{7}\bigg{)}Q+\bigg{(}-\frac{27985893}{4}+8241600\zeta_{3}-\frac{36207\pi^{4}}{2}-1692738\zeta_{5}-\frac{38060\pi^{6}}{7}$
$-2645568\zeta_{3}^{2}+326592\zeta_{7}\bigg{)}Q^{2}+\bigg{(}\frac{1270377}{4}-3338073\zeta_{3}+\frac{64809\pi^{4}}{20}+1459530\zeta_{5}+\frac{1730\pi^{6}}{7}+11124\zeta_{3}^{2}-19278\zeta_{7}\bigg{)}Q^{3}$
$+\bigg{(}\frac{252477}{16}+92178\zeta_{3}+36288\zeta_{5}+9072\zeta_{7}\bigg{)}Q^{4}\bigg{]}\epsilon^{5}~,$
$\Delta_{Q{\tiny\mbox{FP}}}^{5}$ $=$ $\bigg{[}-2852566016+2139424512Q\bigg{]}\epsilon+\bigg{[}\frac{9747487616}{3}-424181248Q-1627608960Q^{2}\bigg{]}\epsilon^{2}+\bigg{[}\frac{77119431952}{9}$ (115)
$+1600458624\zeta_{3}+\bigg{(}-\frac{36719656688}{3}-2589857280\zeta_{3}\bigg{)}Q+\bigg{(}5365027104+905601024\zeta_{3}\bigg{)}Q^{2}-223974720Q^{3}\bigg{]}\epsilon^{3}$
$+\bigg{[}-\frac{18452489398}{3}+16928531152\zeta_{3}+\frac{66685776\pi^{4}}{5}-25646146560\zeta_{5}+\bigg{(}\frac{62653214582}{9}-12498672264\zeta_{3}$
$-21582144\pi^{4}+24847024320\zeta_{5}\bigg{)}Q+\bigg{(}-4161747404-280446528\zeta_{3}+\frac{37733376\pi^{4}}{5}-3130401600\zeta_{5}\bigg{)}Q^{2}$
$+\bigg{(}640256832+94633056\zeta_{3}+135475200\zeta_{5}\bigg{)}Q^{3}-28653588Q^{4}\bigg{]}\epsilon^{4}+\bigg{[}-\frac{12059984664415}{648}-\frac{98689756858\zeta_{3}}{3}$
$+\frac{2116066394\pi^{4}}{15}+141568450614\zeta_{5}-\frac{237464320\pi^{6}}{7}-4542718032\zeta_{3}^{2}-92577398445\zeta_{7}+\bigg{(}\frac{2294461732357}{108}$
$+21292069758\zeta_{3}-\frac{520778011\pi^{4}}{5}-123620853240\zeta_{5}+\frac{230065040\pi^{6}}{7}+7314253776\zeta_{3}^{2}+79133304537\zeta_{7}\bigg{)}Q$
$+\bigg{(}-\frac{27277501615}{6}-237707128\zeta_{3}-\frac{11685272\pi^{4}}{5}+8399505492\zeta_{5}-\frac{28985200\pi^{6}}{7}-2829860928\zeta_{3}^{2}$
$-1796643828\zeta_{7}\bigg{)}Q^{2}+\bigg{(}-215265924-1832415606\zeta_{3}+\frac{3943044\pi^{4}}{5}+1094776020\zeta_{5}+179200\pi^{6}+151863552\zeta_{3}^{2}$
$-230521788\zeta_{7}\bigg{)}Q^{3}+\bigg{(}56025189+32423328\zeta_{3}+10158048\zeta_{5}+10360224\zeta_{7}\bigg{)}Q^{4}-2457216Q^{5}\bigg{]}\epsilon^{5}~,$